### v2

# Exploring the Zero-Shot Capabilities of LLMs Handling Multiple Problems at once

Abstract
Recent studies have proposed placing multiple problems in a single prompt to improve input token utilization for a more efficient LLM inference. We call this multi-problem prompting (MPP), in contrast to conventional single-problem prompting (SPP) that prompts an LLM with a single problem at a time. While MPP has been shown to work comparably well or even better than SPP under few-shot settings, its zero-shot performance is underexplored, which better reveals the innate multiple problem handling capabilities of LLMs. To address that, we study the zero-shot MPP performance of various LLMs on 6 classification and 12 reasoning benchmarks and confirm that LLMs are competent zero-shot multi-problem solvers. We also examine the conditions of effectiveness of zero-shot MPP and explore several model-level factors that may enable MPP. We observe that LLMs consistently perform worse with selecting indices of texts of a given class label and with multiple mixed-source reasoning problems. This indicates a lack of true understanding. We also find that instruction tuning is an important factor than enhances MPP.
[…]
Exploring the Zero-Shot Capabilities of LLMs
Handling Multiple Problems at once
Zhengxiang Wang  and Jordan Kodner  and Owen Rambow Department of Linguistics & Institute for Advanced Computational Science Stony Brook University, Stony Brook, NY, USA {first.last}@stonybrook.edu
1 Introduction
Large language models (LLMs) have recently shown remarkable natural language understanding and reasoning capabilities (OpenAI, 2023; Gemini-Team, 2023; Anthropic, 2024), both under zero-shot (Wei et al., 2022; Kojima et al., 2022) and few-shot (Brown et al., 2020; Wei et al., 2023) settings. Given the non-trivial cost of deploying LLMs at large scale, substantial efforts have been made to make LLM inferences more cost efficient, e.g., distilling a smaller but highly capable model from a large one Abdin et al. (2024); OpenAI (2024).
Besides model-level distillation, recent studies Cheng et al. (2023); Lin et al. (2024); Son et al. (2024) have proposed various prompt-level approaches that place multiple problems in a single prompt, which we collectively call multi-problem prompting (MPP). As illustrated in Fig 1, MPP (e.g., BatchClf) inserts multiple problems $P=(P_{1},P_{2},...,P_{n})$ after a shared context $C$ (e.g., task instruction and/or exemplars) to avoid repeating $C$ $n$ times for each problem as in standard single-problem prompting (SPP, e.g., SingleClf), which
Figure 1: Accuracy of 7 LLMs on 4 task types constructed from SST-2 with up to 5 problems, including standard zero-shot prompting (SingleClf), zero-shot multi-problem prompting (BatchClf), and its reformulations (SelectOne and SelectAll). The text boxes provide simplified task illustrations. See Section 3 for details.
improves input token utilization. In practice, MPP is proposed in few-shot settings, where the context, which contains exemplars, is typically significantly longer than individual problems and an increased inference efficiency is guaranteed for the less repetitions of the shared context Cheng et al. (2023). Empirically, MPP is also found to perform comparably well or even better than SPP across a wide range of domains under few-shot settings Cheng et al. (2023); Lin et al. (2024); Son et al. (2024).
Despite progress in few-shot MPP, the performance of LLMs in zero-shot MPP remains underexplored. This is significant on at least two levels. First, on a practical level, many real-world tasks, such as classification, are typically approached in zero-shot settings Ziems et al. (2024). Moreover, designing few-shot exemplars can be tedious and costly to obtain in practice Kojima et al. (2022); Yasunaga et al. (2024). Second, from a scientific perspective, zero-shot MPP provides deeper insights into the innate capabilities of LLMs concurrently handling multiple tasks, a key indicator of intelligence Meyer and Kieras (1997).
To this end, we evaluate LLMs from 4 model families with zero-shot MPP on multi-problem prompts constructed from 6 classification and 12 reasoning benchmarks. To further understand the conditions of effectiveness of MPP, we also perform MPP with two task reformulations (i.e., SelectOne/SelectAll in Fig 1) and with multiple mixed-source problems leveraging these classification and reasoning benchmarks, respectively. While MPP is meant to be an inference-efficient prompting technique, the main motivation of the paper is to provide scientific understandings of LLM capabilities related to the concurrent handling of multiple problems through comprehensive and systematic contrastive evaluation. For this reason, we also explore additional evidence of strong multiple problem handling capabilities of LLMs and possible model-level factors that may enable MPP.
Our main contributions are as follows:
  • We demonstrate the general effectiveness of zero-shot MPP, when the problems are homogeneous or from a single source.
  • We identify two general conditions under which all tested LLMs perform significantly worse than expected and explore the roles of several model-level factors in MPP.
  • We design a new benchmark ZeMPEB (Zero-shot Multi-Problem Evaluation Benchmark), comprising 53,100 zero-shot multi-problem prompts containing classification and reasoning problems, for evaluating the zero-shot multiple problem handling capabilities of LLMs. ZeMPEB will be released upon acceptance.
The paper proceeds as follows. We reviews related works in Section 2. We perform zero-shot MPP and related experiments with classification and reasoning problems in Sections 3 and 4, respectively. We separate the classification and reasoning experiments in two sections due to the different natures of classification and reasoning, which we leverage for more nuanced evaluations. We explore several model-level factors that may enable MPP in Section 5. We conclude in Section 6.
2 Related Work
We note that current LLM evaluation has predominantly focused on LLM’s performance on single-problem prompts. Each of such prompts presents a single problem and expects one specific answer to that problem, which may implicitly require multi-hop reasoning or multi-step task solving.
Recently, Cheng et al. (2023) propose few-shot MPP named batch prompting that prompts LLMs with problems batched together from single sources following a few batches of equally sized exemplars. They found that few-shot MPP greatly increases LLM inference efficiency while retaining downstream performance with a small batch size (e.g., <6). To ensure that batch prompting works with large batch sizes, Lin et al. (2024) introduce a sampling optimization method that takes a majority vote over repeated permutations of batch samples.
Instead of solving multiple separate problems, Son et al. (2024) prompt LLMs with exemplars to solve multiple related tasks based on a shared problem setup by placing an explicit instruction for each task. They find that instructing LLMs to solve all the tasks at once outperforms solving the individual tasks one by one or in a batch.
The related studies surveyed so far imply that LLMs have strong multiple problem handling capabilities under few-shot settings. To the best of our knowledge, Laskar et al. (2023) is the only study that shows instruction-tuned GPTs can handle 5 short questions sampled from two open-domain QA benchmarks at once zero-shot, but they do not conduct no further experiments. Overall, it is poorly understood how well LLMs can concurrently handle multiple problems zero-shot.
3 MPP: Classification
We first start our MPP experiments with classification problems, followed by experiments with reasoning problems in the next section.
Problem Type Input/Output Format Benchmark # Examples Objective
Classification Single-text input SST-2 (Socher et al., 2013) 1,821 Sentiment analysis
CoLA Warstadt et al. (2019) 1,043 Grammatical acceptability
AGNews (Gulli, 2004) 1,000 Topic classification
Text-pair input MRPC (Dolan and Brockett, 2005) 1,725 Paraphrase detection
SNLI (Bowman et al., 2015) 1,000 Natural language inference
WiC (Pilehvar and Camacho-Collados, 2019) 1,400 Word sense disambiguation
Reasoning Yes/no output StrategyQA (Geva et al., 2021) 2,288 Commonsense reasoning
Coin Flips (Wei et al., 2023) 500 Symbolic reasoning
Multi-choice output AQuA (Ling et al., 2017) 254 Arithmetic reasoning
CommonsenseQA (Talmor et al., 2019) 1,221 Commonsense reasoning
Object tracking (Srivastava et al., 2023) 750 Commonsense reasoning
Bigbench date (Srivastava et al., 2023) 363 Commonsense reasoning
Free-response output Last Letters (Wei et al., 2023) 500 Symbolic reasoning
SVAMP (Patel et al., 2021) 1,000 Arithmetic reasoning
GSM8K (Roy and Roth, 2015) 1,319 Arithmetic reasoning
MultiArith (Patel et al., 2021) 600 Arithmetic reasoning
AddSub (Hosseini et al., 2014) 395 Arithmetic reasoning
SingleEq (Koncel-Kedziorski et al., 2015) 508 Arithmetic reasoning
Table 1: Benchmarks used in Section 3 (Classification) and Section 4 (Reasoning). We use the test splits wherever possible, except for CoLA, StrategyQA, and CommonsenseQA, for which we use the dev splits, since the test splits are not publicly available. For AGNews and SNLI, we randomly sample 1,000 examples from the test splits.
3.1 Data
We consider the following 6 classification benchmarks, as described and referenced in Table 1: SST-2, CoLA, AGNews, MRPC, SNLI, and WiC. These benchmarks are commonly used for NLP evaluation, with SST-2, CoLA, and MRPC appearing in GLUE Wang et al. (2019) and WiC in SuperGLUE (Sarlin et al., 2020). They cover (1) two classification paradigms, i.e., single-text and text-pair classification, and (2) six distinct task objectives.
3.2 Evaluation Tasks
For each benchmark with $m$ class labels $L=(L_{1},...,L_{m})$, we perform standard zero-shot single-problem prompting (SPP) on all the problems to establish baselines for LLMs. We randomly sample 100 distinct sets of 100 problems $\{{P^{(i)}}=(P_{1}^{(i)},P_{2}^{(i)},...,P_{100}^{(i)})\}_{i=1}^{100}$ out of each benchmark and take the first $n$ problems from each problem set $P^{(i)}$ to generate a multi-problem prompt of task size $n$, with the task size for SPP being 1. We consider the following 5 task sizes and ensure that each task size has 100 distinct problem sets: 5/10/20/50/100 for single-text benchmarks and 3/5/10/20/50 for text-pair benchmarks. Note that by design a larger task size only differs from a smaller one by having additional problems given a benchmark, which helps us isolate the effect of task size. In total, this results in 13,500 prompts for all MPP tasks.
As illustrated in Fig 1, there are a total of 4 classification-related task types. These task types are designed to minimally differ from one another such that model performance on a task pair with minimal differences immediately reveals its (lack of) ability in certain aspects we intend to test. We provide further descriptions of the 4 task types below, where every two neighboring task types make a minimally different task pair.
Single Classification (SingleClf)
This is SPP for classification. The prompt indicates the task objective as well as the possible class labels, followed by one target classification problem to be classified.
Batch Classification (BatchClf)
This is MPP for classification. The prompt for BatchClf is nearly identical to that of SingleClf to maximize comparability, expect that it incorporates multiple target classification problems to be classified at once.
Index Selection One Label (SelectOne)
SelectOne is a reformulation of BatchClf. Instead of doing batch classification, SelectOne instructs LLMs to select all indices out of $n$ problems $P$ belonging to each class label $L_{i}$ one at a time in $m$ separate prompts, i.e., $\{j:label(P_{j})=L_{i},j\in[1,n]\}$ for $i\in[1,m]$. The expected output for each $L_{i}$ is “None,” “All,” or a list of applicable indices.
Index Selection All Labels (SelectAll)
SelectAll prompts a system to do SelectOne for all possible class labels $L$ all at once. The expected output is thus a concatenation of SelectOne output for $L$, i.e., $\{(L_{i},\mathrm{SelectOne}(P,L_{i})):i\in[1,m]\}$.
3.3 Experimental Setups
LLMs and evaluation settings
We evaluate 7 LLMs from 4 model families with greedy decoding: Vicuna (13B, Chiang et al., 2023), Mistral 7B (Jiang et al., 2023), Mixtral 8x7B (Jiang et al., 2024), Llama-3 8B and 70B (Instruct, Meta, 2024), GPT-3.5, and GPT-4 (OpenAI, 2023).
Performance metric
We measure the average per-problem accuracy (PPA) to unify the evaluation across the four task types. PPA, defined in Equation 1, is the average accuracy of classifying $n$ problems in each prompt or, in the case of SelectOne, in each set of directly related prompts targeting different class labels.
$\mathrm{PPA}=\frac{1}{n}\sum_{i=1}^{n}\delta(I(P_{i}),A_{i})$ (1)
$I(P_{i})$ is the inferred LLM-generated answer to the $i$th problem in the input prompt, $A_{i}$ is the ground truth, and $\delta(i,j)=1$ iff $i=j$ and 0 otherwise. For the two index selection tasks, $I(P_{i})$ is determined by considering the LLM’s assignments of indices to all class labels. Other than assigning an index to a wrong class label, there are two more error types. First, LLMs may assign an index with more than one class label, i.e., an contradiction error. Second, LLMs may assign no labels to an index at all, namely, an non-excluded middle error.
To compare performance difference, we use Mann-Whitney U tests for significant testing and Cohen’s $d$ Cohen (1969) for measuring effect size.
Figure 2: Aggregated average accuracy of the 7 LLMs on the 4 types of tasks across various task sizes for each benchmark. See Appendix B for the full results for each LLM.
Figure 3: Cumulative error density across positions averaging results from the benchmarks. Task size 3/100: the 3 text-pair/single-text benchmarks; otherwise: all the 6 classification benchmarks. See Appendix B for full results.
Figure 4: Pairwise comparisons of SelectOne and SelectAll for each LLM across different task sizes averaging over results from the 6 benchmarks.
BatchClf vs
SelectOne
BatchClf vs
SelectAll
SelectOne vs
SelectAll
Mean Acc Dif 32.0 12.1 -19.9
Std Dev 16.9 15.3 12.0
Cohen’s $d$ 1.8 0.8 -1.0
Table 2: Pairwise accuracy differences (x vs y = x - y). All the differences are statistically significant and with a large effect size (| Cohen’s $d$ | $\geq 0.8$).
3.4 Results
Overall, we observe that although LLMs are competent zero-shot multi-problem solvers for classification (Section 3.4.1), their performance decreases significantly when the same sets of multiple problems are presented differently (Section 3.4.2).
3.4.1 Zero-shot MPP: SingleClf vs. BatchClf
In line with previous studies Cheng et al. (2023); Lin et al. (2024) on few-shot MPP, we find that zero-shot MPP also performs comparably well or even better than zero-shot SPP in a more cost-efficient way. We further validate this finding by comparing the errors LLMs make under SPP and MPP.
LLMs can handle multiple classifications at once under zero-shot with minimal performance loss.
Although the BatchClf accuracy generally declines as the task size increases (Fig 2), all 7 LLMs achieve accuracy of at least 90% that of SingleClf across the benchmarks most of the time (see Table 5 in Appendix B). Overall, the SingleClf accuracy for the 7 LLMs on the 6 benchmarks is 75.5% and the BatchClf accuracy is 72.3%, a minor 3.2% absolute drop from the former.
Zero-shot MPP can be cost-efficient.
With a general task instruction being shared, zero-shot MPP avoids repeating this instruction for each problem to be solved. Given the strong zero-shot MPP performance, this makes MPP a promising cost-efficient zero-shot prompting technique. To illustrate, we choose for each model/task combination the largest BatchClf task size that achieves at least 95% SingleClf accuracy for that pair. We observe that MPP reduces substantial inference costs for all LLMs run on the 6 benchmarks, ranging from from 30.7% to 82.0% (see Fig 7 in Appendix B).
LLMs make similar errors evenly across positions under multiple problems.
Given the results above, two natural questions arise: do LLMs make similar errors under MPP and how do their errors distribute across positions? For each one of the 180 LLM/benchmark/task size combinations, we use chi-squared tests to compare the proportional error distribution across class labels under SingeClf and BatchClf and compute the cumulative error density across positions under BatchClf. We observe that (1) only in 9 out of 180 (or 5%) cases, error labels are distributed significantly differently between SingleClf and BatchClf (p < 0.05); and (2) surprisingly, LLMs typically do not display a clear positional bias or a serial position effect as known in psychology Murdock (1962), when solving sufficiently many problems at once (Fig 3). This is in contrast to previous studies based on single-problem prompts where LLMs are found to be better at using information from the beginning (primacy bias) or the end (receny bias) of the prompt Liu et al. (2024); Levy et al. (2024).
We hypothesize that the differences in observations may be due to different ways of measuring positional biases. Without a precise definition of positional biases and clear notions of what counts as “beginning,” “middle,” and “end,” we find it necessary to look at cumulative error density (Fig 3) to identify the overall positional error distribution. We leave the arguments for future studies.
Taken together, the fact that LLMs make similar label prediction errors and the lack of an obvious positional bias imply that LLMs can use information equally well across different positions under multiple classification problems. This may explain their strong multiple problem handling capabilities.
3.4.2 MPP Reformulations: Selection Tasks
We further compare LLM’s performance between BatchClf and the two selection tasks.
LLMs perform significantly worse on the selection tasks.
In our experiments, LLMs nearly always perform much better in BatchClf than in SelectOne and SelectAll under the same conditions with a consistent and stable margin, even when the task size is just 3 or 5 (Figs 12). The overall discrepancy in accuracy between BatchClf and the two tasks is large and statistically significant (32% for SelectOne and 10% for SelectAll, see Table 2) and generally increases with a larger task size (Fig 2).
The sharp drop in accuracy may not be human-like, because arguably, humans should at least be able to classify and select a small number (e.g., 3/5) of texts equally well simply by thinking over the problems (i.e., zero-shot). This suggests either LLMs’ lack of true understanding of the problems presented in various forms or their struggle at index selection.
LLMs lack internalized planning and consistently select inconsistently.
Surprisingly, such a consistent and rather stable performance gap also exists between SelectOne and SelectAll in favor of the latter, largely independently of the task size (Fig 2). On average, the SelectOne accuracy is 19.9% lower than the SelectAll accuracy, also with a large and significant effect size (Table 2).
We investigate the reasons in Fig 4, which shows that when asked to select text indices for one class label at a time independently, LLMs almost always assign an index to multiple labels (i.e., contradiction) and leave some indices unselected (i.e., non-excluded middles) more often. This showcases a lack of internalized planning with modern zero-shot LLMs, although different LLMs may make these two types of errors in different proportions. In contrast, when LLMs have to select indices for all labels at once, they are less likely to generate directly illogical answers in a single output as their answer to the $(i+1)_{th}$ label is conditioned by their answer to the $i_{th}$ label. This highlights the purely probabilistic aspect of LLM answer generation.
LLMs can do index selection, but fail to combine it with text classification adequately.
To see if LLMs’ significantly poorer performance with selection tasks is due to their struggle at index selection, we perform a follow-up study. More concretely, we simplify the two index selection tasks by directly replacing each line of text with its gold class label in the prompts for AGNews, which has most labels. We then ask the LLMs to select indices of lines containing each label with minimal modifications to the original task instructions (see Appendix E).
Figure 5: Average accuracy of the 7 LLMs on the two simplified index selection tasks based on AGNews. We provide the original results from BatchClf, SelectOne, and SelectAll for easy comparisons.
Fig 5 shows that LLMs can perform in the simplified selection tasks much better than the original ones, with Llama-3 70B, GPT-3.5, and GPT-4 even achieving (nearly) 100% accuracy across task sizes. Their overall performance in the simplified selection tasks even surpasses their overall BatchClf accuracy. The previously observed performance gap between SelectOne and SelectAll also nearly disappears (see Appendix B for full results). In general, LLMs can indeed do index selection.
In conclusion, it appears that the tested LLMs perform less well on SelectOne and SelectAll as compared to BatchClf because they cannot adequately combine the index selection task and the classification tasks in response to a zero-shot prompt. [^1] We also tried 1-shot-CoT to explicitly prompt LLMs to do BatchClf first and then select indices from there. Although that helps, this basic finding remains. See Appendix B.6. Put differently, the LLMs lack true understanding of the problems presented in different forms, even when the number of problems is quite small (e.g., 3 or 5), which may not be human-like.
4 MPP: Reasoning
We provide the MPP experimental details and results for reasoning problems in this section.
4.1 Data
We further experiment with another 12 reasoning benchmarks (listed and referenced in Table 1), commonly used in LLM evaluation (Kojima et al., 2022; Wei et al., 2023; Zhang et al., 2023, inter alia). These benchmarks test symbolic reasoning (Coin Flips & Last Letters), commonsense reasoning (StrategyQA, CommonsenseQA, Object tracking, & Bigbench date), and arithmetic reasoning (AQuA, SVAMP, GSM8K, MultiArith, AddSub, & SingleEq), and require three answer formats (Yes/No, multiple choice, and free-response).
4.2 Evaluation Tasks
We first test on all the reasoning problems in each benchmark to establish LLM SPP baselines. Two MPP evaluation task types are designed, i.e., single-source and mixed-source multi-problem reasoning, or MultiReasonSS and MultiReasonMS. For both task types, we simply bundle multiple reasoning problems together with an indexed prefix “Qi: ,” where i is the index of each problem starting from “Q1.” We use two headers, namely, “Questions,” followed by the bundled questions, and “Answers,” followed by the answers to be generated by LLMs.
Besides vanilla zero-shot prompting, we also perform zero-shot-CoT following Kojima et al. (2022). [^2] In our early experiments, we found that zero-shot-CoT did not lead to different responses for the classification-related MPP tasks probably due to their novelty, so it was not used. Inspired by Cheng et al. (2023) as well as to control for the number of prompts generated, we consider smaller task sizes from 2 to 10. To ensure a reliable evaluation (e.g., sufficient parsable outputs), we increase the number of prompts from 100 to 300 for each benchmark/task size combination. This results in 21,600 and 18,000 prompts for MultiReasonSS and MultiReasonMS, respectively.
MultiReasonSS
For each one of the 12 benchmarks, we first sample 300 distinct sets of 10 problems. We then use the first 2, 5, and 10 problems from each problem set to make prompts.
MultiReasonMS
To robustly examine an LLM’s performance on mixed-source prompts, we create 6 distinct benchmark combinations based on the 12 benchmarks, each consisting of 10 different benchmarks. For each combination, we randomly sample 300 distinct sets of 10 problems, each coming from one benchmark and with the order of benchmarks randomized across the 300 problem sets. We then take out 2/4/6/8/10 problems from the first 2/4/6/8/10 benchmarks for each benchmark combination across each problem set to make mixed-source prompts. See Appendix B for details.
4.3 Experimental Setups
LLMs and evaluation settings
For budget reasons, we only use two LLMs, i.e., GPT-3.5 and Llama-3 70B. Since Llama-3 models tend to produce reasoning steps even when not instructed to do so, we only prompt GPT-3.5 with zero-shot-CoT. We always prompt with greedy decoding.
Performance metric
We use the per-problem accuracy (PPA), defined in Equation 1, to measure the model performance. For mixed-source prompts containing $k$ problems evenly sampled from $k$ benchmarks, we compute the expected PPA by averaging over the observed SPP performance for each one of these benchmarks.
4.4 Results
We observe that although LLMs can be competent zero-shot multi-problem solvers for reasoning, their performance becomes consistently worse than expected under multiple mixed-source problems. Similar to our arguments in Section 3, the consistent performance declines even with a small number (e.g., 2 or 4) of mixed-source problems may indicate a lack of human-like understanding, as LLMs’ reasoning capabilities are easily impacted by the mixing of problems from different sources.
LLMs can do MultiReasonSS on par with their SPP performance.
Similar to Cheng et al. (2023), we observe in Fig 6 (A) that both LLMs, to varying extents, can handle multiple single-source reasoning problems as well as or even better than when they handle these problems individually, although their MPP performance typically goes down with a larger task size.
When the reasoning problems are from mixed sources, LLMs perform worse than expected.
Interestingly, as shown in Fig 6 (B), the observed MultiReasonMS performance is almost always lower than the expected one computed by averaging over the SPP performance over each benchmark for both LLMs, with and without CoT. Out of 540 model (including GPT-3.5 with zero-shot-CoT) and benchmark pairs, there are only 18.3% cases in which the observed performance is better than the expected one for a given model/benchmark pair by a small margin (mean/std: 2.9%/2.5%). However, for the rest 81.7% cases when the expected performance is better, the margin is larger (mean/std: 13.5%/12.7%). In other words, LLMs typically perform worse in each benchmark when handling multiple reasoning problems from mixed sources.
Figure 6: Average accuracy of GPT-3.5 and Llama-3 70B on multiple single-source (MultiReasonSS) and mixed-source (MultiReasonMS) reasoning problems. (A) MultiReasonSS results on 6 selected benchmarks for space reasons. The other 6 benchmarks show similar results (see Appendix B). (B) MultiReasonMS results across 6 distinct benchmark combinations, where each benchmark contributes a problem to each mixed-source prompt.
Benefits of zero-shot-CoT prompting are transferrable under MPP.
Analogous to the finding that zero-shot-CoT improves LLMs’ reasoning performance under SPP Kojima et al. (2022), GPT-3.5 generally performs better with CoT than without it under zero-shot MPP for both MultiReasonSS and MultiReasonMS. The transferrability of zero-shot-CoT [^3] Similarly, Cheng et al. (2023) show that the benefits of few-shot-CoT are transferrable under MPP. indicates that LLMs can apply CoT over each problem in the prompt and benefit from the generated reasoning steps when solving each problem. This again implies the strong capabilities of LLMs to utilize information across positions under MPP.
5 Exploring Model-level Factors that may Enable MPP
Since so far we have only tested decoder-only and instruction-tuned LLMs, which all show strong performance under MPP, we explore if these two factors enable MPP. For these reasons, we test with greedy decoding (1) three FLAN-T5 models Chung et al. (2022), i.e., Large, XL, XXL; and (2) three pretrained LLMs, i.e., Llama-3 8B (Base, Meta, 2024), GPT-3 1.3B, and GPT-3 175B Brown et al. (2020). We run these 6 LLMs on CoLA at task sizes 1 and 5 under zero-shot settings with results shown in Table 3. Experiments with Coin Flips show similar results, but the LLM outputs are less meaningful, as discussed in Appendix D.
We make three observations from Table 3. First, instruction tuning helps, since pretraned decoder-only LLMs either cannot handle multiple problems at once or their performance is much worse than their instruction-tuned counterparts. However, unlike what Laskar et al. (2023) suggests, instruction tuning may not be a necessary condition for MPP, as both Llama-3 8B and GPT-3 175B can perform reasonably well in BatchClf on CoLA.
Second, FLAN-T5 can barely respond to MPP, regardless of model sizes. We suspect that this is may not be due to their encoder-decoder structures, but other factors such as training data and reinforcement learning from human feedback Christiano et al. (2017). We leave it for future investigation.
Third, scaling model size seems helpful. With other factors being identical, larger models appear to perform better than the smaller ones under MPP.
Overall, instruction tuning appears to be the most important factor that enhances MPP. We leave more careful explorations to future research.
SingleClf BatchClf Avg # Answers
Llama-3 8B (Instruct) 80.5 79.4 5.0
GPT-3.5 84.2 79.6 5.0
Llama-3 8B (Base) 78.5 60.6 5.04
GPT-3 1.3B 63.0 0.0 0.03
GPT-3 175B 66.6 64.4 5.08
FLAN-T5-Large (0.78B) 76.0 NA 1.0
FLAN-T5-XL (3B) 80.2 NA 1.0
FLAN-T5-XXL (11B) 78.2 4.0 1.2
Table 3: SingleClf and BatchClf (task size 5) accuracy (%) of three pretrained LLMs and three FLAN-T5 models on CoLA. We also include the results of Llama-3 8B (Instruct) and GPT-3.5 (likely base model: GPT-3 175B) from Section 3 to compare with their respective base models. The last column is the average number of LLM-generated answers for BatchClf (expects 5). When it is 1, accuracy is not calculated to avoid overinterpretation.
6 Conclusion
In this study, we present a comprehensive evaluation of LLMs with zero-shot MPP. We evaluate various LLMs from 4 model families on single-source multi-problem prompts constructed from 6 classification and 12 reasoning benchmarks. In line with previous few-shot results, we confirm that LLMs are competent multi-problem solvers for classification and reasoning under zero-shot settings. Moreover, we find multiple pieces of evidence that validate the strong innate multiple problem handling capabilities of LLMs, such as the similar classification errors LLMs make under SPP and MPP, the lack of obvious positional biases, and the transferrability of zero-shot-CoT under MPP.
Two conditions are identified under which LLMs show consistent performance declines with MPP: (1) reformulating Batch Classification as index selection tasks; and (2) mixing reasoning problems from different sources in a multi-problem prompt. Noticeably, these performance declines happen even when the number of problems included is rather small (e.g., $\leq$ 5), which may not be human-like and indicates a lack of true understanding. In addition, we explore several model-level factors that may enable MPP and find instruction tuning to be an important factor that enhances MPP.
As a result of our study, we create a new benchmark comprising 53,100 zero-shot multi-problem prompts. We call it ZeMPEB, which stands for Zero-shot Multi-Problem Evaluation Benchmark. We will release ZeMPEB to aid future research.
Limitations
While we have done our best to make our experiments as comprehensive as possible, there is always more work that can be done in a large comparison study, including the addition of even more benchmarks and language models. We were limited by cost, time, and space and have attempted to select an informative and representative sample of experiments. Since we used pre-existing benchmarking datasets, we inherit any errors that they main contain. Finally, despite our efforts to compare different prompts, as is the case with all prompt-based LLM studies, we cannot guarantee that slight differences in the prompts would not meaningfully alter the results.
Ethical Concerns
To the best of our knowledge, all results published in this paper are accurate. All data sources are free, publicly available, and cited in the article. No sensitive data was used which could violate individuals’ privacy or confidentiality.
Appendix A LLM Details
We use a total of 13 LLMs in our study, from Section 3 to Section 5. Table 4 describes the specific versions for these LLMs and highlights their differences in terms of architecture, open weights, Supervised Fine-Tuning (SFT, Wei et al., 2022), and Reinforment Learning from Human Feedback (RLHF, Christiano et al., 2017) etc.
Model Version Architecture Open Weights SFT RLHF MoE # Parameter Context Length
Vicuna (Chiang et al., 2023) v1.5 decoder-only 13B 4,096
Mistral 7B (Jiang et al., 2023) Instruct-v0.2 decoder-only 7B 8,192
Mixtral 8x7B (Jiang et al., 2024) Instruct-v0.1 decoder-only 47B 8,192
Llama-3 8B (Meta, 2024) Instruct decoder-only 8B 8,192
Llama-3 70B (Meta, 2024) Instruct decoder-only 70B 8,192
GPT-3.5 turbo-0125 decoder-only Unknown 16,385
GPT-4 (OpenAI, 2023) turbo-2024-04-09 decoder-only Unknown 128,000
GPT-3 1.3B (Brown et al., 2020) babbage-002 decoder-only 1.3B 16,384
GPT-3 175B (Brown et al., 2020) davinci-002 decoder-only 175B 16,384
Llama-3 8B (Meta, 2024) Base decoder-only 8B 8,192
FLAN-T5 (Chung et al., 2022) Large encoder-decoder 0.78B 512
FLAN-T5 (Chung et al., 2022) XL encoder-decoder 3B 512
FLAN-T5 (Chung et al., 2022) XXL encoder-decoder 11B 512
Table 4: Details about the 13 LLMs used in the study. For Mixtral 8x7B, a Mixture of Experts (MoE) LLM, although each token has access to 47B parameters, but only uses 13B active parameters during inference.
> 90% SCAcc > 80% SCAcc > 75% SCAcc
Vicuna 13B 79.3 93.1 93.1
Mistral 7B 76.7 83.3 100.0
Mixtral 8x7B 63.3 83.3 86.7
Llama-3 8B 73.3 90.0 100.0
Llama-3 70B 80.0 100.0 100.0
GPT-3.5 56.7 83.3 90.0
GPT-4 100.0 100.0 100.0
Overall 75.6 90.4 95.7
Table 5: Percent of time that BatchClf performance surpasses a threshold percent of SingleClf accuracy (SCAcc) across benchmarks.
Appendix B MPP: Classification
This section contains additional details for Section 3. The prompt details for the experiments can be found in Appendix E.
B.1 Full Results
The full results obtained from Section 3 are visualized in Fig 8. We exclude the results of Vicuna on AGNews when task size is 100 because the prompts exceed the model’s context length.
B.2 SingleClf vs. BatchClf
Table 5 indicates the proportion of BatchClf tasks for which each LLM surpasses a threshold percent of corresponding SingleClf performance.
B.3 Zero-shot MPP can be cost-efficient
Fig 7 shows the cost saving rate for each model/task pair at the largest BatchClf task size that achieves at least 95% SingleClf accuracy for that pair.
Figure 7: Cost saving rate (%) per classification based on our experiments. The cost is estimated by both the input and output token counts (using the respective tokenizers), weighted according to the pricing policy from OpenAI and TogetherAI (for non-GPT LLMs) websites.
B.4 Positional Errors under BatchClf
Fig 9 shows the full results regarding the positional errors 7 LLMs make across benchmarks and task sizes. We note that (1) distribution of the positional errors becomes more random (or even) as the task size increases for all LLMs; (2) in most cases, the positional errors distribute nearly randomly, showing no evidence of obvious positional biases, if any; (3) some LLMs may display more severe positional biases on some benchmarks with a certain task size, such as GPT-3.5 on SST-2 with task size 50, but overall this is rare.
B.5 Simplified Index Selection Tasks
Fig 10 shows the full results for the two simplified index selection tasks, along with results for the two original selection tasks and BatchClf, based on AGNews. Clearly, (1) all LLMs perform much better in the simplified tasks than the original ones; (2) Mixtral 8X7B, Llama-3 70B, GPT-3.5, and GPT-4 can do these simplified selection tasks even better than their BatchClf performance consistently, with the latter three achieving nearly 100% accuracy across task sizes; (3) the task complexity hierarchy between SelectOne and SelectAll nearly disappears for all LLMs except Vicua, implying that it may be challenging for Vicuna to perform index selection tasks in general.
Figure 8: Full average accuracy of 7 LLMs on the 4 classification-related tasks across the 6 classification benchmarks.
Figure 9: Cumulative error density across positions for all benchmarks and LLMs across different task sizes. The task size “M/N” on the left side of the plots denotes the task size for the 3 text-pair benchmarks (i.e., MRPC, SNLI, and WiC) and for the 3 single-text benchmarks (i.e., SST-2, CoLA, AGNews), respectively.
Figure 10: Full results for the two simplified index selection tasks versus the original tasks based on AGNews.
Figure 11: Average accuracy of GPT-3.5 and Llama-3 70B on the other 6 reasoning benchmarks with multiple single-source problems. We leave out results where the number of parsable outputs is less than 50, e.g., GPT-3.5 on StrategyQA at task size > 2.
Task SelectOne + CoT SelectAll + CoT BatchClf
Model
Vicuna 23.0 25.8 53.2 57.6 68.7
Mistral 7B 38.3 47.5 56.8 60.5 71.2
Mixtral 8x7B 47.4 38.7 65.1 58.4 73.4
Llama-3 8B 39.0 41.8 62.4 59.0 73.3
Llama-3 70B 59.4 67.1 72.9 79.2 79.4
GPT-3.5 45.5 47.5 66.7 66.3 71.9
GPT-4 66.3 71.8 78.8 81.8 81.9
Overall 45.5 48.6 65.1 66.1 74.3
Table 6: Aggregate average accuracy of SelectOne and SelectAll with and without 1-shot-CoT for each LLM. BatchClf performance is also provided for comparisons.
B.6 CoT Helps the Index Selection Tasks Slightly but the Gap Remains
We use 1-shot-CoT to prompt LLMs to do BatchClf first and then perform the two index selection tasks on the 6 benchmarks with a fixed task size 10. Table 6 shows that while overall LLMs benefit from CoT for both SelectOne and SelectAll, the benefits are generally larger for SelectOne than SelectAll (3.1% versus 1.0% improvement) and not consistent across models with Mixtral 8x7B performing even worse with CoT. The full results provided in Table D also show mixed effects across benchmarks. Interestingly, however, the task complexity hierarchy among BatchClf, SelectOne, and SelectAll remains in virtually all cases. We leave to future studies the investigation of how CoT may further help with the two selection tasks.
Appendix C MPP: Reasoning
This section provides additional details for Section 4. The prompt details for the experiments can be found in Appendix E.
C.1 Construction of Mixed-Source Prompts
We create 6 distinct benchmark combinations based on the 12 benchmarks, each consisting of 10 different benchmarks. When creating these 6 benchmark combinations, we implement the following 2 rules: (1) the first 2 benchmarks must be different across the 6 combinations to cover the 12 benchmarks; (2) the first 2 benchmarks cannot come from SVAMP, GSM8K, MultiArith, AddSub, and SingleEq to maximize the differences between them.
C.2 More Single-Source Results
Fig 11 shows MultiReasonSS results on the other 6 reasoning benchmarks not presented in Fig 6.
# P = 1 # P = 2 Avg # Answers
Llama-3 8B (Instruct) 46.8 50.0 2.1
GPT-3.5 44.8 55.4 2.0
Llama-3 8B (Base) 45.9 (20.4) 33.7 2.7
GPT-3 1.3B 49.0 45.0 6.3
GPT-3 175B 50.0 (43.4) 28.4 6.6
FLAN-T5-Large (0.78B) 46.6 NA 1.0
FLAN-T5-XL (3B) 49.4 NA 1.0
FLAN-T5-XXL (11B) 57.2 NA 1.0
Table 7: Accuracy (%) of three pretrained LLMs and three FLAN-T5 models on Coin Flips with 1 and 2 problems. “# P”: number of problems. When computing accuracy, we treat an LLM response with more than 2 answers as a wrong answer. The numbers in the two parentheses are accuracies when treating “uncertain” answers as wrong answers, instead of discarding them.
Appendix D Additional Results for Section 5
This section describes the results of the three pretrained base LLMs and three FLAN-T5 models on Coin Flips at task sizes 1 and 2 from Section 5, shown in Table 7. Similar to Table 3, Table 7 shows that instruction-tuned models perform much better on multi-problem prompts and that FLAN-T5 models can barely handle multi-problem prompts.
However, after manual inspection, we find that the outputs of the three pretrained models, are often not very sensical with repetitions of the prompts (either partially or entirely). In particular, as shown in Table 7, the two GPT-3 models tend to produce more answers than needed (we set max output tokens to be 200). The answer can also be non-sensical even when well formatted, such as the example output from GPT-3 175B below. Llama-3 8B (Base) overall produce more meaningful outputs, but similar problems also apply. We thus discourage over-interpreting the results in Table 7.
Q1: Yes. The coin is still heads up. Q2: No. The coin is now tails up. Explanation Q1: The coin is still heads up because the coin is not flipped. The coin is flipped only when someone flips the coin. If no one flips the coin, the coin is not flipped. Q2: The coin is now tails up because the coin is flipped. The coin is flipped only when someone flips the coin. If no one flips the coin, the coin is not flipped. Discussion This is a very interesting question. It is a question about the meaning of “flip”. The meaning of “flip” is not clear. It is not clear whether “flip” means “reverse” or “flip”. It is not clear whether "flip" means "flip" or "flip"…{The last sentence repeated for three more times.}…
Benchmark SST-2 CoLA AGNews MRPC SNLI WiC
Model Task
Vicuna SelectOne 38.5 22.2 16.8 15.8 9.9 34.5
+ CoT 52.6 53.9 12.1 12.1 11.5 12.5
SelectAll 66.7 65.7 42.2 67.4 28.2 48.8
+ CoT 81.7 69.2 40.9 63.9 40.9 48.9
BatchClf 89.1 73.2 70.4 66.7 59.3 53.4
[…]
+ CoT 66.5 55.1 44.8 59.4 29.5 29.5
SelectAll 68.3 63.6 48.2 58.4 54.7 47.9
+ CoT 74.4 68.2 57.0 68.5 47.6 47.3
BatchClf 79.6 76.8 78.1 74.0 69.1 49.6
[…]
+ CoT 41.5 67.3 12.9 48.4 25.1 37.2
SelectAll 85.3 68.4 62.9 65.9 56.2 52.0
+ CoT 49.0 67.4 57.8 66.5 58.9 51.1
BatchClf 80.2 80.9 71.6 70.8 73.3 63.7
[…]
+ CoT 85.0 48.9 22.1 50.7 22.3 22.0
SelectAll 85.0 62.9 59.6 64.8 50.1 52.2
+ CoT 85.8 66.5 45.3 64.5 39.8 52.3
BatchClf 92.0 76.8 79.5 62.8 69.2 59.5
[…]
+ CoT 94.7 79.7 75.2 67.9 69.4 15.8
SelectAll 87.3 79.6 76.2 71.7 68.6 53.9
+ CoT 96.1 83.7 85.5 73.9 77.3 58.4
BatchClf 96.4 82.8 86.8 73.5 75.2 61.7
[…]
+ CoT 80.2 71.7 57.9 26.8 26.5 22.1
SelectAll 83.1 79.7 68.3 61.5 55.8 51.6
+ CoT 89.5 77.5 58.8 70.2 50.5 51.2
BatchClf 95.0 81.8 86.4 52.7 57.4 58.0
[…]
+ CoT 84.5 81.9 67.8 60.1 77.0 59.6
SelectAll 92.4 84.0 79.2 73.5 80.8 62.8
+ CoT 91.3 83.7 85.4 76.2 83.3 71.0
BatchClf 96.0 84.0 86.7 72.6 83.8 68.4
Overall SelectOne 68.6 58.0 48.7 37.7 27.6 32.7
+ CoT 72.1 65.5 41.8 46.5 37.3 28.4
SelectAll 81.2 72.0 62.4 66.2 56.3 52.7
+ CoT 81.1 73.7 61.5 69.1 56.9 54.3
BatchClf 89.8 79.5 79.9 67.6 69.6 59.2
Table 8: Full 1-shot-CoT results for all benchmarks and LLMS. The task size is fixed at 10.
Appendix E Full Prompts
E.1 Prompt Templates for SingleClf, BatchClf, SelectOne, and SelectAll
Tables 13 to 14 show the complete prompt templates for the four task types (i.e., SingleClf, BatchClf, SelectOne, and SelectAll) tailored for SST-2, CoLA, AGNews, MPRC, SNLI, and WiC, respectively. While there are differences in the exact wording of a prompt template for each task type across the 6 classification benchmarks, each prompt template type shares a similar underlying structure and can be easily applied to other classification benchmarks.
Throughout our research project, we have also tried prompts with different wordings and structures until we finally unified the prompt designs presented above. For example, we initially asked LLMs to directly generate indices line by line instead of a JSON output for SelectOne and we did not provide any formatted example for SelectAll. We also put the output format instruction in the end of each prompt for SelectAll, instead of in the beginning. Although we observed certain task-level performance variations, which are expected, the overall complexity among the 4 task types (SelectOne > SelectAll > BatchClf > SingeClf) remains unchanged, despite the variations in the prompts. This indicates the overall limited effects of rewording and restructuring prompts.
E.2 Prompt Templates for the Two Index Selection Tasks
Table 15 presents the two prompt templates used for the two simplified index selection tasks based on AGNews.
E.3 One-Shot-CoT Prompt Templates for the Two Index Selection Tasks
To construct a 1-shot-CoT exemplar, we first randomly sampled 10 examples from each benchmark that do not appear in our test examples. We then put these 10 examples into the prompt template for SelectOne or SelectAll, followed by a constructed answer. The answer first explicitly classifies the 10 examples into the corresponding class labels, and then perform index selection based on the generated class labels.
Table 16 shows the complete answer templates for SelectOne and SelectAll for each classification benchmark.
E.4 Prompt Template for Multi-problem Prompts for Reasoning Problems
The prompt template for a multi-problem prompt made up of reasoning problems is straightforward, as described in Section 4. Below is a simple example prompt made up of 2 reasoning problems from CommonsenseQA.
Questions Q1: The person wasn’t bothered by the weather, she had remembered to bring her what? Answer Choices: (A) read book (B) own house (C) apartment (D) more rice (E) warm coat Q2: After working on the car, what did it end up doing? Answer Choices: (A) going too fast (B) last several years (C) honk the horn (D) go fast (E) start running Answers
To enable zero-shot-CoT, we simply append the string “Let’s think step by step.” Kojima et al. (2022) to a zero-shot prompt like the one shown above in a newline (after “Answers”).
Task Prompt template
SingleClf Indicate the sentiment for the following line of text. The sentiment shall be either ‘Positive’ or ‘Negative.’
Text: \$text
Sentiment:
BatchClf Indicate the sentiment for each of the \$num following lines of text. The sentiment shall be either ‘Positive’ or ‘Negative.’
Texts, one per line:
\$texts
The sentiments for each of the \$num lines of text, one per line:
SelectOne Go over the \$num lines of text below and list the index numbers of the lines with \$polarity sentiment according to the following instructions: If none of the texts show \$polarity sentiment, write ‘None.’ If all the texts show \$polarity sentiment, write ‘All.’ Otherwise, provide the index numbers for each text with \$polarity sentiment.
Output your responses in JSON format with the key ‘\$polarity’. A formatted example output is provided below.
{‘\$polarity’: [None/All or index numbers for the texts with \$polarity sentiment]}
Texts, one per line:
\$texts
JSON output:
SelectAll Go over the \$num lines of text below. First, list the index numbers of the lines with positive sentiment. Then, list the index numbers of the lines with negative sentiment. If none of the texts show a particular sentiment, write ‘None.’ If all the texts show a particular sentiment, write ‘All.’ Otherwise, provide the index numbers of the texts that fit a particular category.
Output your responses in JSON format with two keys: ‘positive’ and ‘negative.’ A formatted example output is provided below.
{‘positive’: [None/All or index numbers of positive sentences], ‘negative’: [None/All or index numbers of negative sentences]}
Texts, one per line:
\$texts
JSON output:
Table 9: Prompt templates for SST-2. Words immediately preceded by the dollar sign \$ are placeholders. For the single-text classification task (SST-2, CoLA, AGNews), the sequence of texts in the place of ‘\$texts’ are indexed starting with ‘1’ and each text is separated by a newline.
Task Prompt template
SingleClf Indicate the grammatical acceptability for the following line of text. The acceptability shall be either ‘Acceptable’ or ‘Unacceptable.’
Text: \$text
Grammatical acceptability:
BatchClf Indicate the grammatical acceptabilities for each of the \$num following lines of text. The acceptability shall be either ‘Acceptable’ or ‘Unacceptable.’
Texts, one per line:
\$texts
Grammatical acceptabilities for each of the \$num lines of text, one per line:
SelectOne Go over the \$num lines of text below and list the index numbers of the lines that are grammatically \$acceptability according to the following instructions: If none of the texts are grammatically \$acceptability, write ‘None.’ If all the texts are grammatically \$acceptability, write ‘All.’ Otherwise, provide the index numbers for each grammatically \$acceptability text.
Output your responses in JSON format with the key ‘\$acceptability’. A formatted example output is provided below.
{‘\$acceptability’: [None/All or index numbers of \$acceptability sentences]}
Texts, one per line:
\$texts
JSON output:
SelectAll Go over the \$num lines of text below. First, list the index numbers of the lines that are grammatically acceptable. Then, list the index numbers of the lines that are grammatically unacceptable. If none of the sentences show a particular acceptability, write ‘None.’ If all the sentences show a particular acceptability, write ‘All.’ Otherwise, provide the index numbers of the texts that fit a particular category.
Output your responses in JSON format with two keys ‘acceptable’ and ‘unacceptable.’ A formatted example output is provided below.
{‘acceptable’: [None/All or index numbers of acceptable texts], ‘unacceptable’: [None/All or index numbers of unacceptable texts]}
Texts, one per line:
\$texts
JSON output:
Table 10: Prompt templates for CoLA.
Task Prompt template
SingleClf Classify which news category the following line of text belongs to among the following four categories: ‘Business,’ ‘Sports,’ ‘World,’ and ‘Sci/Tech.’
Text: \$text
News category:
BatchClf Classify which news category each of the \$num following lines of text belongs to among the following four categories: ‘Business,’ ‘Sports,’ ‘World,’ and ‘Sci/Tech.’
Texts, one per line:
\$texts
News categories for each of the \$num lines of text, one per line:
SelectOne This is a news classification task in which each line of text belongs to one of four categories ‘Business,’ ‘Sports,’ ‘World,’ and ‘Sci/Tech.’
Go over the \$num lines of text below and list the index numbers of the lines that can be classified as \$category according to the following instructions: If none of the texts can be classified as \$category, write ‘None.’ If all the texts can be classified as \$category, write ‘All.’ Otherwise, provide the index numbers of the texts that can be classified as \$category.
Output your responses in JSON format with the key ‘\$category’. A formatted example output is provided below.
{‘\$category’: [None/All or index numbers of the texts that can be classified as \$category]}
Texts, one per line:
\$texts
JSON output:
SelectAll This is a news classification task in which each line of text belongs to one of four categories ‘Business,’ ‘Sports,’ ‘World,’ and ‘Sci/Tech.’
Go over the \$num lines of text below and list the index numbers of the lines that belong to each category according to the following instructions: If none of the texts can be classified as a particular category, write ‘None.’ If all the texts can be classified as a particular category, write ‘All.’ Otherwise, provide the index numbers of the texts that can be classified as the category.
Output your responses in JSON format with the following keys: ‘business,’ ‘sports,’ ‘world,’ and ‘sci/tech.’ A formatted example output is provided below.
{‘business’: [None/All or index numbers of texts in ‘business’ category], ‘sports’: [None/All or index numbers of texts in ‘sports’ category], ‘world’: [None/All or index numbers of texts in ‘world’ category], ‘sci/tech’: [None/All or index numbers of texts in sci/tech category]}
Texts, one per line:
\$texts
JSON output:
Table 11: Prompt templates for AGNews.
Task Prompt template
SingleClf Compare text A with text B and determine if text A is a paraphrase of text B. Respond with ‘Yes’ if text A is a paraphrase, and ‘No’ if it is not.
\$text
Answer:
BatchClf Compare text A with text B for the following \$num text pairs and determine if text A is a paraphrase of text B line by line. Respond with ‘Yes’ if text A is a paraphrase, and ‘No’ if it is not. Provide your answers line by line.
\$texts
Answers:
SelectOne Go over the \$num text pairs below and list the index numbers of the text pairs where text A \$be a paraphrase of text B according to the following instructions: If none of the text pairs satisfy this condition, write ‘None.’ If all the text pairs satisfy this condition, write ‘All.’ Otherwise, provide the index numbers of the text pairs where text A \$be a paraphrase of text B.
Output your responses in JSON format with the key ‘answer’. A formatted example output is provided below.
{‘answer’: [None/All or index numbers of the text pairs where text A \$be a paraphrase of text B]}
Here are the text pairs:
\$texts
JSON output:
SelectAll Go over the \$num text pairs below. First, list the index numbers of the text pairs that contain paraphrases. Then, list the index numbers of the text pairs that contain non-paraphrases. If none of the text pairs satisfy a condition, write ‘None.’ If all the text pairs satisfy a condition, write ‘All.’ Otherwise, provide the index numbers of the text pairs that satisfy each condition.
Output your responses in JSON format with two keys: ‘yes’ for paraphrases and ‘no’ for non-paraphrases. A formatted example output is provided below.
{‘yes’: [None/All or index numbers of text pairs that contain paraphrases], ‘no’: [None/All or index numbers of text pairs that contain non-paraphrases]}
Here are the text pairs:
\$texts
JSON output:
Table 12: Prompt templates for MRPC. For the text-pair classification task (MRPC, SNLI, WiC), the sequence of text pairs in the place of ‘\$texts’ are indexed starting with ‘1’ and each text pair is separated by two newlines (each text pair ends with a newline be design, followed by another newline before the next text pair).
Task Prompt template
SingleClf Given the following premise and hypothesis, determine the inference relation between them. Respond with ‘Entailment’ if the hypothesis logically follows from the premise, ‘Contradiction’ if they are in direct opposition, and ‘Neutral’ if neither applies.
\$text
Inference relation:
BatchClf Given the following \$num pairs of premises and hypotheses, determine the inference relation for each pair line by line. Respond with ‘Entailment’ if the hypothesis entails the premise, and ‘Contradiction’ if they contradict. If neither is the case, respond with ‘Neutral.’ Provide your answers line by line.
\$texts
Inference relations for the \$num text pairs provided above:
SelectOne Go over the \$num text pairs below and list the index numbers of the text pairs where the inference relation between the premise and the hypothesis is \$relationship according to the following instructions: If none of the text pairs contain \$relationship inference relation, write ‘None.’ If all text pairs contain \$relationship inference relation, write ‘All.’ Otherwise, provide the index numbers of the text pairs where the inference relation between the premise and the hypothesis is \$relationship.
Output your responses in JSON format with the key ‘\$relationship’. A formatted example output is provided below.
‘\$relationship’: [None/All or index numbers of text pairs that contain \$relationship inference relation]
Here are the text pairs:
\$texts
JSON output:
SelectAll Go over the \$num text pairs below. First, list the index numbers of the text pairs that contain entailment inference relation. Then, select all text pairs that contain contradiction inference relation. Finally, select all text pairs that contain neutral inference relation. If none of the text pairs satisfy a condition, write ‘None.’ If all the text pairs belong satisfy a condition, write ‘All.’ Otherwise, provide the index numbers of the text pairs that satisfy each condition.
Output your responses in JSON format with three keys: ‘entailment’, ‘contradiction’, and ‘neutral’. A formatted example output is provided below.
{‘entailment’: [None/All or index numbers of text pairs that contain entailment inference relation], ‘contradiction’: [None/All or index numbers of text pairs that contain contradiction inference relation], ‘neutral’: [None/All or index numbers of text pairs that contain neutral inference relation]}
Here are the text pairs:
\$texts
JSON output:
Table 13: Prompt templates for SNLI.
Task Prompt template
SingleClf Analyze the usage of the given target word in the two subsequent contexts. The target word may appear in various grammatical forms in each context. Respond with ‘Yes’ if it maintains the same meaning across both contexts, and ‘No’ if it does not.
\$text
Answer:
BatchClf Analyze the usage of the following \$num target words in the two contexts that immediately follow them. These target words may appear in different grammatical forms across the two subsequent contexts. Determine if each target word maintains the same meaning in the two subsequent contexts. Provide your answers line by line, indicating ‘Yes’ if it does and ‘No’ if it does not.
\$texts
Answers:
SelectOne Analyze the following \$num target words and determine the index numbers of the target words where the same meaning \$be maintained across the two contexts that immediately follow them. These target words may appear in different grammatical forms in each context. If none of the target words satisfy this condition, write ‘None.’. If all the target words satisfy this condition, write ‘All.’ Otherwise, provide the index numbers.
Output your responses in JSON format with the key ‘answer’. A formatted example output is provided below.
{‘answer’: [None/All or index numbers of the target words where the same meaning \$be maintained in the two subsequent contexts]}
Here are the target words along with their contexts:
\$texts
JSON output:
SelectAll Analyze the following \$num target words, which may appear in different grammatical forms in the two subsequent contexts. First, list the index numbers of target words that maintain the same meaning in the two subsequent contexts. Then, list the index numbers of target words that do not maintain the same meaning in the two subsequent contexts. If none of the target words satisfy a condition, write ‘None.’ If all the target words satisfy a condition, write ‘All.’ Otherwise, provide the index numbers of the target words that satisfy each condition.
Output your responses in JSON format with two keys: ‘yes’ for target words used with consistent meanings and ‘no’ for those used with inconsistent meanings. A formatted example output is provided below.
{‘yes’: [None/All or index numbers of target words used with consistent meanings], ‘no’: [None/All or index numbers of target words used with inconsistent meanings]}
Here are the target words along with their contexts:
\$texts
JSON output:
Table 14: Prompt templates for WiC.
Task Prompt template
SelectOne Simplified Go over the \$num lines of text below and list the index numbers of the lines that contain the word ‘\$category’ according to the following instructions: If none of the texts contain the word ‘\$category,’ write ‘None.’ If all the texts contain the word ‘\$category,’ write ‘All.’ Otherwise, provide the index numbers of the texts that contain the word ‘\$category’ each on a separate line.
Texts, one per line:
\$texts
‘None,’ ‘All,’ or the index numbers of the texts that contain the word ‘\$category,’ one per line:
SelectAll Simplified In this task, each line of text contains one of four words ’Business,’ ’Sports,’ ’World,’ and ’Sci/Tech.’
Go over the \$num lines of text below and list the index numbers of the lines that contain each word according to the following instructions: If none of the texts contain a particular word, write ‘None.’ If all the texts contain a particular word, write ‘All.’ Otherwise, provide the index numbers of the texts that contain each word.
Output your responses in JSON format with the following keys: ‘business,’ ‘sports,’ ‘world,’ and ‘sci/tech.’ A formatted example output is provided below.
{‘business’: [None/All or index numbers of texts containing ‘Business’], ‘sports’: [None/All or index numbers of texts containing ‘Sports’], ‘world’: [None/All or index numbers of texts containing ‘World’], ‘sci/tech’: [None/All or index numbers of texts containing ‘Sci/Tech’]}
Texts, one per line:
\$texts
JSON output:
Table 15: Prompt templates for simplified SelectOne and SelectAll based on AGNews. For SelectOne Simplified, we used an earlier prompt template that asks LLMs to produce outputs line by line, instead of a JSON object. As illustrated earlier, we find the effect of the two output formats for SelectOne to be minimal. Therefore, having a JSON output should not make a meaningful difference.
Benchmark Task Answer template
SST-2 SelectOne To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
From here, we see that lines {indices} show positive sentiment. Therefore, the answer in JSON format is as follows: {final answer}
SST-2 SelectAll To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
From here, we see that lines {indices} show positive sentiment, and lines {indices} show negative sentiment. Therefore, the answer in JSON format is as follows: {final answer}
CoLA SelectOne To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
From here, we see that texts in lines {indices} are unacceptable. Therefore, the answer in JSON format is as follows: {final answer}
CoLA SelectAll To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
From here, we see that texts in lines {indices} are acceptable and texts in lines {indices} are unacceptable. Therefore, the answer in JSON format is as follows: {final answer}
AGNews SelectOne To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
From here, we see that texts in lines {indices} can be classified as ‘Sports.’. Therefore, the answer in JSON format is as follows: {final answer}
AGNews SelectAll To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
From here, we see that texts in lines {indices} can be classified as ‘Business,’ texts in lines {indices} can be classified as ‘Sports,’ texts in lines {indices} can be classified as ‘World,’ and texts in lines {indices} are ‘Sci/Tech.’ Therefore, the answer in JSON format is as follows: {final answer}
MRPC SelectOne To solve this task, let’s first determine if text A is a paraphrase of text B for the 10 lines of text above, one per line: {labels}
From there, we see that text pairs in lines {indices} are paraphrases. Therefore, the answer in JSON format is as follows: {final answer}
MRPC SelectAll To solve this task, let’s first determine if text A is a paraphrase of text B for the 10 lines of text above, one per line: {labels}
From there, we see that text pairs in lines {indices} are paraphrases and text pairs in lines {indices} are not. Therefore, the answer in JSON format is as follows: {final answer}
SNLI SelectOne To solve this task, let’s first determine the inference relation between the premise and the hypothesis for the 10 lines of text above, one per line: {labels}
From there, we see that text pairs in lines {indices} contain entailment inference relation. Therefore, the answer in JSON format is as follows: {final answer}
SNLI SelectAll To solve this task, let’s first determine the inference relation between the premise and the hypothesis for the 10 lines of text above, one per line: {labels}
From there, we see that text pairs in lines {indices} contain entailment inference relation, text pairs in lines {indices} contain contradiction inference relation, and text pairs in lines {indices} contain neutral inference relation. Therefore, the answer in JSON format is as follows: {final answer}
WiC SelectOne To solve this task, let’s first determine if the target word is used with consistent meanings in the two subsequent contexts for the num lines of text above, one per line: {labels}
From there, we see that text pairs in lines {indices} contain use the target words with inconsistent meanings. Therefore, the answer in JSON format is as follows: {final answer}
WiC SelectAll To solve this task, let’s first determine if the target word is used with consistent meanings in the two subsequent contexts for the num lines of text above, one per line: {labels}
From there, we see that text pairs in lines {indices} contain use the target words with inconsistent meanings and text pairs in lines {indices} do not. Therefore, the answer in JSON format is as follows: {final answer}
Table 16: Answer templates for the two selection tasks for each classification benchmark. {indices}: a list of indices seperated by comma. {final answer}: answer in a JSON format specified by the SelectOne or SelectAll prompt.

### v3

@@
-Exploring the Zero-Shot Capabilities of LLMs Handling Multiple Problems at once
+Evaluating LLMs with Multiple Problems at once
 Abstract
-Recent studies have proposed placing multiple problems in a single prompt to improve input token utilization for a more efficient LLM inference.
-We call this multi-problem prompting (MPP), in contrast to conventional single-problem prompting (SPP) that prompts an LLM with a single problem at a time.
-While MPP has been shown to work comparably well or even better than SPP under few-shot settings, its zero-shot performance is underexplored, which better reveals the innate multiple problem handling capabilities of LLMs.
-To address that, we study the zero-shot MPP performance of various LLMs on 6 classification and 12 reasoning benchmarks and confirm that LLMs are competent zero-shot multi-problem solvers.
-We also examine the conditions of effectiveness of zero-shot MPP and explore several model-level factors that may enable MPP.
-We observe that LLMs consistently perform worse with selecting indices of texts of a given class label and with multiple mixed-source reasoning problems.
-This indicates a lack of true understanding.
-We also find that instruction tuning is an important factor than enhances MPP.
+This paper shows the benefits and fruitfulness of evaluating LLMs with multiple problems at once, a paradigm we call multi-problem evaluation (MPE).
+Unlike conventional single-problem evaluation, where a prompt presents a single problem and expects one specific answer, MPE places multiple problems together in a single prompt and assesses how well an LLM answers all these problems in a single output.
+Leveraging 6 classification and 12 reasoning benchmarks that already exist, we introduce a new benchmark called ZeMPE (Zero-shot Multi-Problem Evaluation), comprising 53,100 zero-shot multi-problem prompts.
+We experiment with a total of 13 LLMs from 5 model families on ZeMPE to present a comprehensive and systematic MPE.
+Our results show that LLMs are capable of handling multiple problems from a single data source as well as handling them separately, but there are conditions this multiple problem handling capability falls short.
+In addition, we perform in-depth further analyses and explore model-level factors that may enable multiple problem handling capabilities in LLMs.
+We release our corpus and code [^1] https://github.com/jaaack-wang/multi-problem-eval-llm to facilitate future research.
 […]
 […]
-Exploring the Zero-Shot Capabilities of LLMs
-Handling Multiple Problems at once
+Evaluating LLMs with Multiple Problems at once
 Zhengxiang Wang  and Jordan Kodner  and Owen Rambow Department of Linguistics & Institute for Advanced Computational Science Stony Brook University, Stony Brook, NY, USA {first.last}@stonybrook.edu
 1 Introduction
-Large language models (LLMs) have recently shown remarkable natural language understanding and reasoning capabilities (OpenAI, 2023; Gemini-Team, 2023; Anthropic, 2024), both under zero-shot (Wei et al., 2022; Kojima et al., 2022) and few-shot (Brown et al., 2020; Wei et al., 2023) settings.
-Given the non-trivial cost of deploying LLMs at large scale, substantial efforts have been made to make LLM inferences more cost efficient, e.g., distilling a smaller but highly capable model from a large one Abdin et al.
-(2024); OpenAI (2024).
-Besides model-level distillation, recent studies Cheng et al.
+Thanks to the advances in both GPU hardware and algorithms (Dai et al., 2019; Beltagy et al., 2020; Dao et al., 2022; Ding et al., 2024; Chen et al., 2024, inter alia), large language models (LLMs) have been developed with increasingly larger context windows (e.g., 8K, 128K, 2M).
+To leverage the extended context windows, recent studies Cheng et al.
 (2023); Lin et al.
 (2024); Son et al.
-(2024) have proposed various prompt-level approaches that place multiple problems in a single prompt, which we collectively call multi-problem prompting (MPP).
-As illustrated in Fig 1, MPP (e.g., BatchClf) inserts multiple problems $P=(P_{1},P_{2},...,P_{n})$ after a shared context $C$ (e.g., task instruction and/or exemplars) to avoid repeating $C$ $n$ times for each problem as in standard single-problem prompting (SPP, e.g., SingleClf), which
-Figure 1: Accuracy of 7 LLMs on 4 task types constructed from SST-2 with up to 5 problems, including standard zero-shot prompting (SingleClf), zero-shot multi-problem prompting (BatchClf), and its reformulations (SelectOne and SelectAll).
-The text boxes provide simplified task illustrations.
-See Section 3 for details.
-improves input token utilization.
-In practice, MPP is proposed in few-shot settings, where the context, which contains exemplars, is typically significantly longer than individual problems and an increased inference efficiency is guaranteed for the less repetitions of the shared context Cheng et al.
+(2024) have proposed various prompting strategies that place multiple problems in a single prompt, which we collectively call multi-problem prompting (MPP).
+The basic idea of MPP is to place multiple problems after a shared context $C$ (e.g., task instruction and/or exemplars) to avoid repeating $C$ for each problem as in standard single-problem prompting (SPP), which improves input token utilization and reduces LLM inference costs per problem.
+In this study, we evaluate a wide range of LLMs with multiple problems at once through MPP, a paradigm we call multi-problem evaluation (MPE) Wang et al.
 (2025).
-Empirically, MPP is also found to perform comparably well or even better than SPP across a wide range of domains under few-shot settings Cheng et al.
-(2023); Lin et al.
-(2024); Son et al.
-(2024).
-Despite progress in few-shot MPP, the performance of LLMs in zero-shot MPP remains underexplored.
-This is significant on at least two levels.
-First, on a practical level, many real-world tasks, such as classification, are typically approached in zero-shot settings Ziems et al.
-(2024).
-Moreover, designing few-shot exemplars can be tedious and costly to obtain in practice Kojima et al.
-(2022); Yasunaga et al.
-(2024).
-Second, from a scientific perspective, zero-shot MPP provides deeper insights into the innate capabilities of LLMs concurrently handling multiple tasks, a key indicator of intelligence Meyer and Kieras (1997).
-To this end, we evaluate LLMs from 4 model families with zero-shot MPP on multi-problem prompts constructed from 6 classification and 12 reasoning benchmarks.
-To further understand the conditions of effectiveness of MPP, we also perform MPP with two task reformulations (i.e., SelectOne/SelectAll in Fig 1) and with multiple mixed-source problems leveraging these classification and reasoning benchmarks, respectively.
-While MPP is meant to be an inference-efficient prompting technique, the main motivation of the paper is to provide scientific understandings of LLM capabilities related to the concurrent handling of multiple problems through comprehensive and systematic contrastive evaluation.
-For this reason, we also explore additional evidence of strong multiple problem handling capabilities of LLMs and possible model-level factors that may enable MPP.
+[^2] While MPE is achieved through MPP, MPP can be used for purposes other than evaluation, e.g., knowledge retrieval, question answering, and other use cases.
+It it thus necessary to distinguish MPP from MPE and SPP from SPE.
+While the main goal of MPP is to improve the cost-efficiency of LLM inference, we view MPE primarily as a valuable evaluation paradigm for probing LLM capabilities, rather than merely a cost-saving engineering strategy.
+Unlike conventional single-problem evaluation that assesses an LLM’s ability to answer a single problem through SPP, MPE assesses an LLM’s ability to concurrently handle multiple problems at once or in a single output.
+Understanding the multiple problem handling capabilities of LLMs is an important research question because it gives us a foundational insight into how LLMs operate over multi-problem inputs that can be sufficiently long and use information from individual problems contained within each multi-problem input.
+To enable a comprehensive and systematic MPE, we introduce ZeMPE (Zero-shot Multi-Problem Evaluation), a new benchmark comprising 53,1000 zero-shot multi-problem prompts.
+ZeMPE is synthetically generated by leveraging 6 classification and 12 reasoning benchmarks that already exist and are widely used.
+Moreover, ZeMPE includes various types of evaluation tasks to allow for deep and nuanced analyses, taking into account how multiple problems are presented in the prompt and whether these problems are sampled from the same data source or not.
+We do not mix classification and reasoning problems together due to the different natures of these two types of problems and to not make our experiments confounding.
 Our main contributions are as follows:
-  • We demonstrate the general effectiveness of zero-shot MPP, when the problems are homogeneous or from a single source.
-  • We identify two general conditions under which all tested LLMs perform significantly worse than expected and explore the roles of several model-level factors in MPP.
-  • We design a new benchmark ZeMPEB (Zero-shot Multi-Problem Evaluation Benchmark), comprising 53,100 zero-shot multi-problem prompts containing classification and reasoning problems, for evaluating the zero-shot multiple problem handling capabilities of LLMs.
-ZeMPEB will be released upon acceptance.
-The paper proceeds as follows.
-We reviews related works in Section 2.
-We perform zero-shot MPP and related experiments with classification and reasoning problems in Sections 3 and 4, respectively.
-We separate the classification and reasoning experiments in two sections due to the different natures of classification and reasoning, which we leverage for more nuanced evaluations.
-We explore several model-level factors that may enable MPP in Section 5.
-We conclude in Section 6.
+  • We show that LLMs are capable of handling multiple classification or reasoning problems from a single data source as well as handling them separately zero-shot.
+We present multiple pieces of evidence, in addition to direct performance comparisons to validate this.
+  • We demonstrate that just like few-shot MPP, zero-shot MPP can be highly cost-efficient.
+  • We identify two general conditions under which LLMs perform significantly worse than expected when presented with multiple problems and explore the roles of several model-level factors that may enable their multiple problem handling capabilities.
+  • We release a new MPE benchmark called ZeMPE to facilitate future MPE studies.
 2 Related Work
@@
 (2023) propose few-shot MPP named batch prompting that prompts LLMs with problems batched together from single sources following a few batches of equally sized exemplars.
-They found that few-shot MPP greatly increases LLM inference efficiency while retaining downstream performance with a small batch size (e.g., <6).
+They find that few-shot MPP greatly increases LLM inference efficiency while retaining downstream performance with a small batch size (e.g., <6).
 To ensure that batch prompting works with large batch sizes, Lin et al.
@@
 They find that instructing LLMs to solve all the tasks at once outperforms solving the individual tasks one by one or in a batch.
-The related studies surveyed so far imply that LLMs have strong multiple problem handling capabilities under few-shot settings.
-To the best of our knowledge, Laskar et al.
-(2023) is the only study that shows instruction-tuned GPTs can handle 5 short questions sampled from two open-domain QA benchmarks at once zero-shot, but they do not conduct no further experiments.
-Overall, it is poorly understood how well LLMs can concurrently handle multiple problems zero-shot.
-3 MPP: Classification
-We first start our MPP experiments with classification problems, followed by experiments with reasoning problems in the next section.
+In addition to these few-shot studies, Laskar et al.
+(2023) shows that instruction-tuned GPTs can handle 5 short questions sampled from two open-domain QA benchmarks at once zero-shot, but the base GPT models can barely perform the task.
+To the best of our knowledge, Wang et al.
+(2025) present the first systematic evaluation of LLMs’ zero-shot ability to tackle multiple homogeneous classification problems drawn from six standard benchmarks.
+They show that, while LLMs can usually solve several such classifications in a single prompt with accuracy comparable to handling them one-by-one, their performance deteriorates sharply when the prompt instead asks them to return the indices of texts belonging to each class–a shortfall that remains consistent across models, prompting conditions, and experimental settings.
+Building on top of Wang et al.
+(2025), this study examines a total of 13 LLMs on 18 existing benchmarks, including 12 reasoning benchmarks that are not part of Wang et al.
+(2025)’s evaluation.
+Besides from reaffirming Wang et al.
+(2025)’s finding that LLMs are capable of handling multiple problems from a single data source as well as handling them separately, we perform in-depth further analyses to both validate such capabilities and expose their limitations.
+Moreover, we explore model-level factors that may enable LLM’s strong multiple problem handling capabilities.
+3 Multi-Problem Evaluation
+This section compares single-problem evaluation (SPE) and multi-problem evaluation (MPE) and introduces ZeMPE, a new MPE benchmark.
+3.1 SPE vs. MPE
+SPE assesses an LLM’s ability to solve a type of problem by prompting the LLM with such a problem one at a time.
+In contrast, MPE places multiple problems together that can be of a same or different types and evaluates how well an LLM handles them.
+A simple example of a multi-problem task would bundle multiple classification or QA problems together and ask LLMs to solve them sequentially.
+3.2 Benefits of MPE
+MPE has at least three advantages over SPE.
+Lesser Data Contamination Concerns
+First, it is less likely for LLMs to encounter exact multi-problem prompts during pre-training because of the combinatory nature of constructing prompts from multiple problems.
+This helps mitigate a growing data contamination concern in modern large-scale pre-training (Jacovi et al., 2023; Sainz et al., 2023).
+Improved Controllability and Interpretability of Evaluation
+Second, since we can manipulate what kind of problems and how many problems to include, we know exactly which problem an LLM gets wrong or right across positions in the prompts.
+This enables us to construct a well controlled and easily interpretable evaluation.
+High Feasibility and Adaptability
+Third, our study demonstrates that leveraging the rich existing benchmarks to create a new multi-problem task is cheap, easy to implement, and highly adaptable.
+The most laborious component is the prompt design, which, once done, can easily be applied to a set of benchmarks with minimal adaptation.
 Problem Type Input/Output Format Benchmark # Examples Objective
@@
 SingleEq (Koncel-Kedziorski et al., 2015) 508 Arithmetic reasoning
-Table 1: Benchmarks used in Section 3 (Classification) and Section 4 (Reasoning).
+Table 1: Existing benchmarks we use to construct ZeMPE.
 We use the test splits wherever possible, except for CoLA, StrategyQA, and CommonsenseQA, for which we use the dev splits, since the test splits are not publicly available.
 For AGNews and SNLI, we randomly sample 1,000 examples from the test splits.
-3.1 Data
-We consider the following 6 classification benchmarks, as described and referenced in Table 1: SST-2, CoLA, AGNews, MRPC, SNLI, and WiC.
-These benchmarks are commonly used for NLP evaluation, with SST-2, CoLA, and MRPC appearing in GLUE Wang et al.
+3.3 ZeMPE
+We describe how we construct ZeMPE as well as how we evaluate LLMs on it.
+3.3.1 Data
+We use 6 classification and 12 reasoning benchmarks, as described and referenced in Table 1, to ensure a comprehensive and systematic evaluation.
+The classification benchmarks are commonly used for NLP evaluation, with SST-2, CoLA, and MRPC appearing in GLUE Wang et al.
 (2019) and WiC in SuperGLUE (Sarlin et al., 2020).
-They cover (1) two classification paradigms, i.e., single-text and text-pair classification, and (2) six distinct task objectives.
-3.2 Evaluation Tasks
-For each benchmark with $m$ class labels $L=(L_{1},...,L_{m})$, we perform standard zero-shot single-problem prompting (SPP) on all the problems to establish baselines for LLMs.
-We randomly sample 100 distinct sets of 100 problems $\{{P^{(i)}}=(P_{1}^{(i)},P_{2}^{(i)},...,P_{100}^{(i)})\}_{i=1}^{100}$ out of each benchmark and take the first $n$ problems from each problem set $P^{(i)}$ to generate a multi-problem prompt of task size $n$, with the task size for SPP being 1.
-We consider the following 5 task sizes and ensure that each task size has 100 distinct problem sets: 5/10/20/50/100 for single-text benchmarks and 3/5/10/20/50 for text-pair benchmarks.
-Note that by design a larger task size only differs from a smaller one by having additional problems given a benchmark, which helps us isolate the effect of task size.
-In total, this results in 13,500 prompts for all MPP tasks.
-As illustrated in Fig 1, there are a total of 4 classification-related task types.
-These task types are designed to minimally differ from one another such that model performance on a task pair with minimal differences immediately reveals its (lack of) ability in certain aspects we intend to test.
-We provide further descriptions of the 4 task types below, where every two neighboring task types make a minimally different task pair.
-Single Classification (SingleClf)
-This is SPP for classification. The prompt indicates the task objective as well as the possible class labels, followed by one target classification problem to be classified.
-Batch Classification (BatchClf)
-This is MPP for classification.
-The prompt for BatchClf is nearly identical to that of SingleClf to maximize comparability, expect that it incorporates multiple target classification problems to be classified at once.
-Index Selection One Label (SelectOne)
-SelectOne is a reformulation of BatchClf.
-Instead of doing batch classification, SelectOne instructs LLMs to select all indices out of $n$ problems $P$ belonging to each class label $L_{i}$ one at a time in $m$ separate prompts, i.e., $\{j:label(P_{j})=L_{i},j\in[1,n]\}$ for $i\in[1,m]$.
-The expected output for each $L_{i}$ is “None,” “All,” or a list of applicable indices.
-Index Selection All Labels (SelectAll)
-SelectAll prompts a system to do SelectOne for all possible class labels $L$ all at once.
-The expected output is thus a concatenation of SelectOne output for $L$, i.e., $\{(L_{i},\mathrm{SelectOne}(P,L_{i})):i\in[1,m]\}$.
-3.3 Experimental Setups
-LLMs and evaluation settings
-We evaluate 7 LLMs from 4 model families with greedy decoding: Vicuna (13B, Chiang et al., 2023), Mistral 7B (Jiang et al., 2023), Mixtral 8x7B (Jiang et al., 2024), Llama-3 8B and 70B (Instruct, Meta, 2024), GPT-3.5, and GPT-4 (OpenAI, 2023).
-Performance metric
-We measure the average per-problem accuracy (PPA) to unify the evaluation across the four task types.
+They cover two classification paradigms (single-text and text-pair) and six distinct task objectives.
+The 12 reasoning benchmarks are widely utilized in LLM evaluation (Kojima et al., 2022; Wei et al., 2023; Zhang et al., 2023, inter alia).
+These benchmarks test symbolic reasoning (Coin Flips & Last Letters), commonsense reasoning (StrategyQA, CommonsenseQA, Object tracking, & Bigbench date), and arithmetic reasoning (AQuA, SVAMP, GSM8K, MultiArith, AddSub, & SingleEq), and require three answer formats (Yes/No, multiple choice, and free-response).
+3.3.2 Evaluation Tasks and Prompt Design
+We separate the classification and reasoning problems, due to their different natures and to avoid confounding experiments, when designing multi-problem evaluation tasks.
+Unlike previous related studies Cheng et al.
+(2023); Lin et al.
+(2024); Son et al.
+(2024) that evaluate LLMs on multi-problem prompts under few-shot settings, our evaluation tasks are all zero-shot, which are rather underexplored, as shown in Section 2.
+Moreover, zero-shot MPE is significant on at least two levels.
+First, on a practical level, many real-world tasks, such as classification, are typically approached in zero-shot settings Ziems et al.
+(2024).
+Moreover, designing few-shot exemplars can be tedious and costly to obtain in practice Kojima et al.
+(2022); Yasunaga et al.
+(2024).
+Second, from a scientific perspective, zero-shot MPP may provide deeper insights into the innate capabilities of LLMs concurrently handling multiple tasks.
+The evaluation tasks are as follows with the full prompt templates for each task in Appendix E.
+Classification-Related Tasks
+We call the standard classification task via SPP Single Classification or SingleClf, which serves as a baseline to be compared with MPE tasks.
+When an LLM is prompted to solve multiple homogeneous classification problems through MPP, this task is known as Batch Classification or BatchClf.
+Index Selection One Label (SelectOne) and Index Selection All Labels (SelectAll) are two reformulations of BatchClf.
+Instead of making multiple classifications under BatchClf, these two tasks instruct LLMs to select indices of text falling into each class label, either independently in m separate prompts (SelectOne) or altogether in a single prompt (SelectAll), where m is the number of class labels in a benchmark.
+We design the two selection tasks to test LLM’s understanding of the classifications performed under BatchClf.
+Since selection tasks of size $n$ may have anywhere from 0 to $n$ correct indices per class, spurious correlations are less likely during our evaluation, given the combinatory answer space.
+For each of the four tasks above, we start by describing the task in the prompt and then include one or multiple classification problems afterwards.
+LLMs are instructed to solve these problems according to the specific task requirements.
+Reasoning-Related Tasks
+We first test on all the reasoning problems in each benchmark to establish LLM SPP baselines.
+Two MPE tasks are designed, i.e., single-source and mixed-source multi-problem reasoning, or MultiReasonSS and MultiReasonMS.
+For both tasks, we bundle multiple reasoning problems together with an indexed prefix “Qi: ,” where i is the index of each problem starting from “Q1.” We use two headers, namely, “Questions,” before the bundled questions, and “Answers,” before answers to be generated by LLMs.
+Unlike prompts designed for classification-related tasks, there is no shared task description, since each question already describes its own unique problem to be solved.
+3.3.3 ZeMPE Composition
+We define task size $n$ as the number of classification or reasoning problems included in a prompt.
+We construct a multi-problem prompt with all problems sampled from the same benchmark, except for MultiReasonMS where we sample one question from each of $k$ reasoning benchmarks to construct an $k$-problem prompt.
+In total, ZeMPE comprises 53,100 zero-shot multi-problem prompts containing classification and reasoning problems.
+Classification-Related Tasks
+For each classification benchmarks, we consider 5 task sizes and ensure that each task size has 100 distinct prompt instances: 5, 10, 20, 50, and 100 for single-text benchmarks and 3, 5, 10, 20, and 50 for text-pair benchmarks.
+To isolate the effect of task size, a larger task size only differs from a smaller one by having additional problems given a benchmark; and to isolate the effect of task, different MPE tasks share the same sets of problems given a task size and a benchmark.
+In total, this results in 13,500 prompts for classification-related MPE tasks.
+Reasoning-Related Tasks
+Besides vanilla zero-shot prompting, we also perform zero-shot-CoT following Kojima et al.
+(2022).
+[^3] In our early experiments, we found that zero-shot-CoT did not lead to different responses for the classification-related MPE tasks probably due to their novelty, so it was not used.
+Inspired by Cheng et al.
+(2023) as well as to control for the number of prompts generated, we consider smaller task sizes from 2 to 10.
+To ensure a reliable evaluation (e.g., sufficient parsable outputs), we increase the number of prompts from 100 to 300 for each benchmark/task size combination.
+For each reasoning benchmark, we consider task sizes 2, 5, and 10 for MultiReasonSS.
+To robustly examine an LLM’s performance on mixed-source prompts, we create 6 distinct benchmark combinations based on the 12 benchmarks, each consisting of 10 different benchmarks.
+For each benchmark combination, we consider the first 2, 4, 6, 8, and 10 benchmarks (also equals the respective task sizes) in the combination for MultiReasonMS.
+We also control the effects of task size and task with careful sampling, similar to what we did above.
+This results in 21,600 and 18,000 prompts for MultiReasonSS and MultiReasonMS, respectively.
+4 Experiments
+This section first describes the experimental setups and then reports and discusses the results.
+4.1 LLMs and Evaluation Settings
+We evaluate 7 LLMs from 4 model families with greedy decoding for the four classification-related tasks: Vicuna (13B, Chiang et al., 2023), Mistral 7B (Jiang et al., 2023), Mixtral 8x7B (Jiang et al., 2024), Llama-3 8B and 70B (Instruct, Meta, 2024), GPT-3.5, and GPT-4 (OpenAI, 2023).
+See Appendix A for the details about the LLMs used.
+Given the consistent results we observed across LLMs with the classification-related tasks and for budget reasons, we only use two LLMs with greedy decoding, i.e., GPT-3.5 and Llama-3 70B.
+Since Llama-3 models tend to produce reasoning steps even when not instructed to do so, we only prompt GPT-3.5 with zero-shot-CoT.
+Figure 1: Average accuracy of the 7 LLMs on the 4 classification-related task across task sizes for each benchmark.
+Figure 2: Cost/Accuracy Ratio (lower is better) for the 7 LLMs on the 6 benchmarks for SingleClf (task size=1) and BatchClf (otherwise).
+We use the average (input + output) token count per classification as the proxy for the actual inference costs, calculated on the basis of the input and output tokens.
+The plot for MRPC is log-scale for the y-axis.
+4.2 Performance Metric
+We measure the average per-problem accuracy (PPA) to unify the evaluation across all proposed tasks.
 PPA, defined in Equation 1, is the average accuracy of classifying $n$ problems in each prompt or, in the case of SelectOne, in each set of directly related prompts targeting different class labels.
 $\mathrm{PPA}=\frac{1}{n}\sum_{i=1}^{n}\delta(I(P_{i}),A_{i})$ (1)
-$I(P_{i})$ is the inferred LLM-generated answer to the $i$th problem in the input prompt, $A_{i}$ is the ground truth, and $\delta(i,j)=1$ iff $i=j$ and 0 otherwise.
-For the two index selection tasks, $I(P_{i})$ is determined by considering the LLM’s assignments of indices to all class labels.
+where $I(P_{i})$ is the inferred LLM-generated answer to the $i$th problem in the input prompt, $A_{i}$ is the ground truth, and $\delta(i,j)=1$ iff $i=j$ and 0 otherwise.
+For SelectOne and SelectAll, $I(P_{i})$ is determined by considering the LLM’s assignments of indices to all class labels.
 Other than assigning an index to a wrong class label, there are two more error types.
-First, LLMs may assign an index with more than one class label, i.e., an contradiction error.
-Second, LLMs may assign no labels to an index at all, namely, an non-excluded middle error.
+First, LLMs may assign an index with more than one class label, i.e., a contradiction error.
+Second, LLMs may assign no labels to an index at all, namely, a non-excluded middle error.
+For MultiReasonMS prompts containing $k$ problems evenly sampled from $k$ benchmarks, we compute the expected PPA by averaging over the observed SPP performance for each benchmark.
 To compare performance difference, we use Mann-Whitney U tests for significant testing and Cohen’s $d$ Cohen (1969) for measuring effect size.
-Figure 2: Aggregated average accuracy of the 7 LLMs on the 4 types of tasks across various task sizes for each benchmark. See Appendix B for the full results for each LLM.
-Figure 3: Cumulative error density across positions averaging results from the benchmarks.
-Task size 3/100: the 3 text-pair/single-text benchmarks; otherwise: all the 6 classification benchmarks.
-See Appendix B for full results.
-Figure 4: Pairwise comparisons of SelectOne and SelectAll for each LLM across different task sizes averaging over results from the 6 benchmarks.
+4.3 Classification-Related Results
+In line with previous studies (Cheng et al., 2023; Lin et al., 2024) on few-shot MPP, we observe that while large language models (LLMs) demonstrate strong zero-shot classification capabilities and prompting with multiple problems can be cost-efficient, their performance degrades significantly when the same sets of problems are presented in a different format.
+LLMs can handle multiple classifications at once under zero-shot with minimal performance loss.
+Although the BatchClf accuracy generally declines as the task size increases (Fig 1), all 7 LLMs achieve accuracy of at least 90% that of SingleClf across the benchmarks most of the time (see Table 5 in Appendix B).
+Overall, the SingleClf accuracy for the 7 LLMs on the 6 benchmarks is 75.5% and the BatchClf accuracy is 72.3%, a minor 3.2% absolute drop from the former.
+Interestingly, for SNLI almost all LLMs perform better in BatchClf than in SingleClf across all the task sizes (3-50) and GPT-4 consistently achieves a BatchClf accuracy near or better than the SingleClf accuracy under all conditions (see Fig 7 in Appendix B).
+Zero-shot MPP can be cost-efficient.
+Single-problem prompting can waste input tokens by redundantly repeating a shared task instruction.
+Multi-problem prompting reduces this redundancy, and this saving is larger the more problems are combined in a single prompt.
+Because performance tends to decline slowly as the number of tasks increases, this yields a favorable cost-accuracy ratio as the number of tasks increases (Fig 2).
+We only encountered two outliers, Vicuna on MRPC with task size 50 where the average input is 3,645 tokens and the context window is 4,096 tokens, and Mixtral 8x7B at $\geq$ 50 on AGNews.
+While it is of course up to downstream users to determine what cost-accuracy is right for them, this is likely beneficial for many use cases where similar prompts are repeated frequently.
+To illustrate, we choose for each model/task combination the largest BatchClf task size that achieves at least 95% SingleClf accuracy for that pair.
+We observe that MPP reduces substantial inference costs for all LLMs run on the 6 benchmarks, ranging from from 30.7% to 82.0% (see Fig 6 in Appendix B).
 BatchClf vs
@@
 Table 2: Pairwise accuracy differences (x vs y = x - y). All the differences are statistically significant and with a large effect size (| Cohen’s $d$ | $\geq 0.8$).
-3.4 Results
-Overall, we observe that although LLMs are competent zero-shot multi-problem solvers for classification (Section 3.4.1), their performance decreases significantly when the same sets of multiple problems are presented differently (Section 3.4.2).
-3.4.1 Zero-shot MPP: SingleClf vs. BatchClf
-In line with previous studies Cheng et al.
-(2023); Lin et al.
-(2024) on few-shot MPP, we find that zero-shot MPP also performs comparably well or even better than zero-shot SPP in a more cost-efficient way.
-We further validate this finding by comparing the errors LLMs make under SPP and MPP.
-LLMs can handle multiple classifications at once under zero-shot with minimal performance loss.
-Although the BatchClf accuracy generally declines as the task size increases (Fig 2), all 7 LLMs achieve accuracy of at least 90% that of SingleClf across the benchmarks most of the time (see Table 5 in Appendix B).
-Overall, the SingleClf accuracy for the 7 LLMs on the 6 benchmarks is 75.5% and the BatchClf accuracy is 72.3%, a minor 3.2% absolute drop from the former.
-Zero-shot MPP can be cost-efficient.
-With a general task instruction being shared, zero-shot MPP avoids repeating this instruction for each problem to be solved.
-Given the strong zero-shot MPP performance, this makes MPP a promising cost-efficient zero-shot prompting technique.
-To illustrate, we choose for each model/task combination the largest BatchClf task size that achieves at least 95% SingleClf accuracy for that pair.
-We observe that MPP reduces substantial inference costs for all LLMs run on the 6 benchmarks, ranging from from 30.7% to 82.0% (see Fig 7 in Appendix B).
-LLMs make similar errors evenly across positions under multiple problems.
-Given the results above, two natural questions arise: do LLMs make similar errors under MPP and how do their errors distribute across positions?
-For each one of the 180 LLM/benchmark/task size combinations, we use chi-squared tests to compare the proportional error distribution across class labels under SingeClf and BatchClf and compute the cumulative error density across positions under BatchClf.
-We observe that (1) only in 9 out of 180 (or 5%) cases, error labels are distributed significantly differently between SingleClf and BatchClf (p < 0.05); and (2) surprisingly, LLMs typically do not display a clear positional bias or a serial position effect as known in psychology Murdock (1962), when solving sufficiently many problems at once (Fig 3).
-This is in contrast to previous studies based on single-problem prompts where LLMs are found to be better at using information from the beginning (primacy bias) or the end (receny bias) of the prompt Liu et al.
-(2024); Levy et al.
-(2024).
-We hypothesize that the differences in observations may be due to different ways of measuring positional biases.
-Without a precise definition of positional biases and clear notions of what counts as “beginning,” “middle,” and “end,” we find it necessary to look at cumulative error density (Fig 3) to identify the overall positional error distribution.
-We leave the arguments for future studies.
-Taken together, the fact that LLMs make similar label prediction errors and the lack of an obvious positional bias imply that LLMs can use information equally well across different positions under multiple classification problems.
-This may explain their strong multiple problem handling capabilities.
-3.4.2 MPP Reformulations: Selection Tasks
-We further compare LLM’s performance between BatchClf and the two selection tasks.
 LLMs perform significantly worse on the selection tasks.
-In our experiments, LLMs nearly always perform much better in BatchClf than in SelectOne and SelectAll under the same conditions with a consistent and stable margin, even when the task size is just 3 or 5 (Figs 12).
+In our experiments, LLMs nearly always perform much better in BatchClf than in SelectOne and SelectAll under the same conditions with a consistent and stable margin, even when the task size is just 3 or 5 (Fig 1).
 The overall discrepancy in accuracy between BatchClf and the two tasks is large and statistically significant (32% for SelectOne and 10% for SelectAll, see Table 2) and generally increases with a larger task size (Fig 1).
 The sharp drop in accuracy may not be human-like, because arguably, humans should at least be able to classify and select a small number (e.g., 3/5) of texts equally well simply by thinking over the problems (i.e., zero-shot).
-This suggests either LLMs’ lack of true understanding of the problems presented in various forms or their struggle at index selection.
-LLMs lack internalized planning and consistently select inconsistently.
 Surprisingly, such a consistent and rather stable performance gap also exists between SelectOne and SelectAll in favor of the latter, largely independently of the task size (Fig 1).
 On average, the SelectOne accuracy is 19.9% lower than the SelectAll accuracy, also with a large and significant effect size (Table 2).
-We investigate the reasons in Fig 4, which shows that when asked to select text indices for one class label at a time independently, LLMs almost always assign an index to multiple labels (i.e., contradiction) and leave some indices unselected (i.e., non-excluded middles) more often.
-This showcases a lack of internalized planning with modern zero-shot LLMs, although different LLMs may make these two types of errors in different proportions.
-In contrast, when LLMs have to select indices for all labels at once, they are less likely to generate directly illogical answers in a single output as their answer to the $(i+1)_{th}$ label is conditioned by their answer to the $i_{th}$ label.
-This highlights the purely probabilistic aspect of LLM answer generation.
-LLMs can do index selection, but fail to combine it with text classification adequately.
-To see if LLMs’ significantly poorer performance with selection tasks is due to their struggle at index selection, we perform a follow-up study.
-More concretely, we simplify the two index selection tasks by directly replacing each line of text with its gold class label in the prompts for AGNews, which has most labels.
-We then ask the LLMs to select indices of lines containing each label with minimal modifications to the original task instructions (see Appendix E).
-Figure 5: Average accuracy of the 7 LLMs on the two simplified index selection tasks based on AGNews. We provide the original results from BatchClf, SelectOne, and SelectAll for easy comparisons.
-Fig 5 shows that LLMs can perform in the simplified selection tasks much better than the original ones, with Llama-3 70B, GPT-3.5, and GPT-4 even achieving (nearly) 100% accuracy across task sizes.
-Their overall performance in the simplified selection tasks even surpasses their overall BatchClf accuracy.
-The previously observed performance gap between SelectOne and SelectAll also nearly disappears (see Appendix B for full results).
-In general, LLMs can indeed do index selection.
-In conclusion, it appears that the tested LLMs perform less well on SelectOne and SelectAll as compared to BatchClf because they cannot adequately combine the index selection task and the classification tasks in response to a zero-shot prompt.
-[^1] We also tried 1-shot-CoT to explicitly prompt LLMs to do BatchClf first and then select indices from there.
-Although that helps, this basic finding remains.
-See Appendix B.6.
-Put differently, the LLMs lack true understanding of the problems presented in different forms, even when the number of problems is quite small (e.g., 3 or 5), which may not be human-like.
-4 MPP: Reasoning
-We provide the MPP experimental details and results for reasoning problems in this section.
-4.1 Data
-We further experiment with another 12 reasoning benchmarks (listed and referenced in Table 1), commonly used in LLM evaluation (Kojima et al., 2022; Wei et al., 2023; Zhang et al., 2023, inter alia).
-These benchmarks test symbolic reasoning (Coin Flips & Last Letters), commonsense reasoning (StrategyQA, CommonsenseQA, Object tracking, & Bigbench date), and arithmetic reasoning (AQuA, SVAMP, GSM8K, MultiArith, AddSub, & SingleEq), and require three answer formats (Yes/No, multiple choice, and free-response).
-4.2 Evaluation Tasks
-We first test on all the reasoning problems in each benchmark to establish LLM SPP baselines.
-Two MPP evaluation task types are designed, i.e., single-source and mixed-source multi-problem reasoning, or MultiReasonSS and MultiReasonMS.
-For both task types, we simply bundle multiple reasoning problems together with an indexed prefix “Qi: ,” where i is the index of each problem starting from “Q1.” We use two headers, namely, “Questions,” followed by the bundled questions, and “Answers,” followed by the answers to be generated by LLMs.
-Besides vanilla zero-shot prompting, we also perform zero-shot-CoT following Kojima et al.
-(2022).
-[^2] In our early experiments, we found that zero-shot-CoT did not lead to different responses for the classification-related MPP tasks probably due to their novelty, so it was not used.
-Inspired by Cheng et al.
-(2023) as well as to control for the number of prompts generated, we consider smaller task sizes from 2 to 10.
-To ensure a reliable evaluation (e.g., sufficient parsable outputs), we increase the number of prompts from 100 to 300 for each benchmark/task size combination.
-This results in 21,600 and 18,000 prompts for MultiReasonSS and MultiReasonMS, respectively.
-MultiReasonSS
-For each one of the 12 benchmarks, we first sample 300 distinct sets of 10 problems. We then use the first 2, 5, and 10 problems from each problem set to make prompts.
-MultiReasonMS
-To robustly examine an LLM’s performance on mixed-source prompts, we create 6 distinct benchmark combinations based on the 12 benchmarks, each consisting of 10 different benchmarks.
-For each combination, we randomly sample 300 distinct sets of 10 problems, each coming from one benchmark and with the order of benchmarks randomized across the 300 problem sets.
-We then take out 2/4/6/8/10 problems from the first 2/4/6/8/10 benchmarks for each benchmark combination across each problem set to make mixed-source prompts.
-See Appendix B for details.
-4.3 Experimental Setups
-LLMs and evaluation settings
-For budget reasons, we only use two LLMs, i.e., GPT-3.5 and Llama-3 70B.
-Since Llama-3 models tend to produce reasoning steps even when not instructed to do so, we only prompt GPT-3.5 with zero-shot-CoT.
-We always prompt with greedy decoding.
-Performance metric
-We use the per-problem accuracy (PPA), defined in Equation 1, to measure the model performance.
-For mixed-source prompts containing $k$ problems evenly sampled from $k$ benchmarks, we compute the expected PPA by averaging over the observed SPP performance for each one of these benchmarks.
-4.4 Results
+4.4 Reasoning-Related Results
 We observe that although LLMs can be competent zero-shot multi-problem solvers for reasoning, their performance becomes consistently worse than expected under multiple mixed-source problems.
-Similar to our arguments in Section 3, the consistent performance declines even with a small number (e.g., 2 or 4) of mixed-source problems may indicate a lack of human-like understanding, as LLMs’ reasoning capabilities are easily impacted by the mixing of problems from different sources.
+Similar to our arguments in last section, the consistent performance declines even with a small number (e.g., 2 or 4) of mixed-source problems may indicate a lack of human-like understanding, as LLMs’ reasoning capabilities are easily impacted by the mixing of problems from different sources.
 LLMs can do MultiReasonSS on par with their SPP performance.
@@
 (A) MultiReasonSS results on 6 selected benchmarks for space reasons.
-The other 6 benchmarks show similar results (see Appendix B).
+The other 6 benchmarks show similar results (see Appendix C).
 (B) MultiReasonMS results across 6 distinct benchmark combinations, where each benchmark contributes a problem to each mixed-source prompt.
@@
 This again implies the strong capabilities of LLMs to utilize information across positions under MPP.
-5 Exploring Model-level Factors that may Enable MPP
+5 Further Analyses
+Figure 4: Cumulative error density across positions averaging results from the benchmarks.
+Task size 3/100: the 3 text-pair/single-text benchmarks; otherwise: all the 6 classification benchmarks.
+See Appendix B for full results.
+Figure 5: Pairwise comparisons of SelectOne and SelectAll for each LLM across different task sizes averaging over results from the 6 benchmarks.
+This section provide further analyses to better understand LLMs’ multiple problem handling capabilities, their limitation, and what enables such capabilities.
+It starts with two error analyses for BatchClf, aiming to see whether models make similar prediction errors and positional errors under BatchClf compared to SingleClf.
+We then investigate the reason why SelectAll appears to be harder than SelectOne, which seems counter-intuitive.
+Lastly, we explore model-level factors that may enable LLMs to receive and handle multiple problems at once.
+5.1 BatchClf Error Analysis
+Given the strong BatchClf results, two natural questions arise: do LLMs make similar errors under MPP and how do their errors distribute across positions?
+For each one of the 180 LLM/benchmark/task size combinations, we use chi-squared tests to compare the proportional error distribution across class labels under SingeClf and BatchClf and compute the cumulative error density across positions under BatchClf.
+We observe that (1) only in 9 out of 180 (or 5%) cases, error labels are distributed significantly differently between SingleClf and BatchClf (p < 0.05); and (2) surprisingly, LLMs typically do not display a clear positional bias or a serial position effect as known in psychology Murdock (1962), when solving sufficiently many problems at once (Fig 4).
+This is in contrast to previous studies based on single-problem prompts where LLMs are found to be better at using information from the beginning (primacy bias) or the end (receny bias) of the prompt Liu et al.
+(2024); Levy et al.
+(2024).
+Taken together, the fact that LLMs make similar label prediction errors and the lack of an obvious positional bias imply that LLMs can use information equally well across different positions under multiple classification problems.
+This may explain their strong multiple problem handling capabilities.
+5.2 Why is SelectAll Harder than SelectOne
+We investigate the reasons in Fig 5, which shows that when asked to select text indices for one class label at a time independently, LLMs almost always assign an index to multiple labels (i.e., contradiction) and leave some indices unselected (i.e., non-excluded middles) more often.
+This showcases a lack of internalized planning with modern zero-shot LLMs, although different LLMs may make these two types of errors in different proportions.
+In contrast, when LLMs have to select indices for all labels at once, they are less likely to generate directly illogical answers in a single output as their answer to the $(i+1)_{th}$ label is conditioned by their answer to the $i_{th}$ label.
+5.3 Exploring Model-level Factors that may Enable MPP
 Since so far we have only tested decoder-only and instruction-tuned LLMs, which all show strong performance under MPP, we explore if these two factors enable MPP.
@@
 We run these 6 LLMs on CoLA at task sizes 1 and 5 under zero-shot settings with results shown in Table 3.
-Experiments with Coin Flips show similar results, but the LLM outputs are less meaningful, as discussed in Appendix D.
-We make three observations from Table 3.
-First, instruction tuning helps, since pretraned decoder-only LLMs either cannot handle multiple problems at once or their performance is much worse than their instruction-tuned counterparts.
+Experiments with Coin Flips show similar results, but the LLM outputs are less meaningful, as discussed in Appendix D.2.
+We make the following observations from Table 3.
+Instruction tuning helps.
+This is because pretrained decoder-only LLMs either cannot handle multiple problems at once or their performance is much worse than their instruction-tuned counterparts.
 However, unlike what Laskar et al.
-(2023) suggests, instruction tuning may not be a necessary condition for MPP, as both Llama-3 8B and GPT-3 175B can perform reasonably well in BatchClf on CoLA.
-Second, FLAN-T5 can barely respond to MPP, regardless of model sizes.
-We suspect that this is may not be due to their encoder-decoder structures, but other factors such as training data and reinforcement learning from human feedback Christiano et al.
-(2017).
+(2023) suggests, instruction tuning may not be a necessary condition for MPP, since both Llama-3 8B and GPT-3 175B can perform reasonably well in BatchClf on CoLA.
+FLAN-T5 can barely respond to MPP, regardless of model sizes.
+We suspect that this may not be due to their encoder-decoder structures, but other factors such as training data and reinforcement learning from human feedback Christiano et al.
+(2017), which FLAN-T5 models lack.
 We leave it for future investigation.
-Third, scaling model size seems helpful. With other factors being identical, larger models appear to perform better than the smaller ones under MPP.
+Scaling model size seems helpful.
+With other factors being identical, larger models appear to perform better than the smaller ones under MPP.
+For example, FLAN-T5-Large is outperformed by both FLAN-T5-XL and -XXL.
+Furthermore, while GPT-3 1.3B and FLAN-T5-XL can perform SingleClf close to or even better than GPT-3 175B and FLAN-T5-XXL, only the larger models can do BatchClf to varying extents–the two smaller models cannot do the task at all.
+Final remark.
 Overall, instruction tuning appears to be the most important factor that enhances MPP. We leave more careful explorations to future research.
@@
 Table 3: SingleClf and BatchClf (task size 5) accuracy (%) of three pretrained LLMs and three FLAN-T5 models on CoLA.
-We also include the results of Llama-3 8B (Instruct) and GPT-3.5 (likely base model: GPT-3 175B) from Section 3 to compare with their respective base models.
+We also include the results of Llama-3 8B (Instruct) and GPT-3.5 (likely base model: GPT-3 175B) from Section 4.3 to compare with their respective base models.
 The last column is the average number of LLM-generated answers for BatchClf (expects 5).
@@
 6 Conclusion
-In this study, we present a comprehensive evaluation of LLMs with zero-shot MPP.
+In this study, we present a comprehensive and systematic MPE of LLMs.
 We evaluate various LLMs from 4 model families on single-source multi-problem prompts constructed from 6 classification and 12 reasoning benchmarks.
@@
 Moreover, we find multiple pieces of evidence that validate the strong innate multiple problem handling capabilities of LLMs, such as the similar classification errors LLMs make under SPP and MPP, the lack of obvious positional biases, and the transferrability of zero-shot-CoT under MPP.
+Leveraging the strong multiple problem handling capabilities, we show that zero-shot MPP can be cost-efficient.
 Two conditions are identified under which LLMs show consistent performance declines with MPP: (1) reformulating Batch Classification as index selection tasks; and (2) mixing reasoning problems from different sources in a multi-problem prompt.
@@
 In addition, we explore several model-level factors that may enable MPP and find instruction tuning to be an important factor that enhances MPP.
+Overall, our experiment demonstrate surprisingly consistent observations across different LLMs and across multi-problem prompts constructed from various benchmarks.
+This consistency indicates the reliability and fruitfulness of MPE as an evaluation paradigm.
 As a result of our study, we create a new benchmark comprising 53,100 zero-shot multi-problem prompts.
-We call it ZeMPEB, which stands for Zero-shot Multi-Problem Evaluation Benchmark.
-We will release ZeMPEB to aid future research.
+We call it ZeMPE, which stands for Zero-shot Multi-Problem Evaluation.
+We release ZeMPE to aid future MPE research.
+Acknowledgements
+We are grateful for the supports from the Institute for Advanced Computational Science (IACS) at Stony Brook University, in particular the free GPT access it provides.
+Zhengxiang Wang is supported by IACS’s Junior Researcher Award since Fall 2024.
+We thank three anonymous reviewers from the Generation, Evaluation & Metrics (GEM) Workshop 2025 for their helpful comments.
+The paper was presented at All Things Language And Computation (ATLAC) organized by the NLP reading group at Stony Brook University.
+We also thank the audience there for their discussions and feedback about the paper.
+Zhengxiang Wang and Owen Rambow were supported in part by funding from the Defense Advanced Research Projects Agency (DARPA) under Contracts No.
+HR01121C0186, No.
+HR001120C0037, and PR No.
+HR0011154158.
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of DARPA.
 Limitations
@@
 Appendix A LLM Details
-We use a total of 13 LLMs in our study, from Section 3 to Section 5.
+We use a total of 13 LLMs in our study.
 Table 4 describes the specific versions for these LLMs and highlights their differences in terms of architecture, open weights, Supervised Fine-Tuning (SFT, Wei et al., 2022), and Reinforment Learning from Human Feedback (RLHF, Christiano et al., 2017) etc.
@@
 Table 5: Percent of time that BatchClf performance surpasses a threshold percent of SingleClf accuracy (SCAcc) across benchmarks.
-Appendix B MPP: Classification
-This section contains additional details for Section 3. The prompt details for the experiments can be found in Appendix E.
+Appendix B Classification-Related Results
+This section contains additional details for Section 4.3. The prompt details for the experiments can be found in Appendix E.
 B.1 Full Results
-The full results obtained from Section 3 are visualized in Fig 8. We exclude the results of Vicuna on AGNews when task size is 100 because the prompts exceed the model’s context length.
+The full results obtained from Section 4.3 are visualized in Fig 7. We exclude the results of Vicuna on AGNews when task size is 100 because the prompts exceed the model’s context length.
 B.2 SingleClf vs. BatchClf
@@
 The cost is estimated by both the input and output token counts (using the respective tokenizers), weighted according to the pricing policy from OpenAI and TogetherAI (for non-GPT LLMs) websites.
-B.4 Positional Errors under BatchClf
-Fig 9 shows the full results regarding the positional errors 7 LLMs make across benchmarks and task sizes.
-We note that (1) distribution of the positional errors becomes more random (or even) as the task size increases for all LLMs; (2) in most cases, the positional errors distribute nearly randomly, showing no evidence of obvious positional biases, if any; (3) some LLMs may display more severe positional biases on some benchmarks with a certain task size, such as GPT-3.5 on SST-2 with task size 50, but overall this is rare.
-B.5 Simplified Index Selection Tasks
-Fig 10 shows the full results for the two simplified index selection tasks, along with results for the two original selection tasks and BatchClf, based on AGNews.
-Clearly, (1) all LLMs perform much better in the simplified tasks than the original ones; (2) Mixtral 8X7B, Llama-3 70B, GPT-3.5, and GPT-4 can do these simplified selection tasks even better than their BatchClf performance consistently, with the latter three achieving nearly 100% accuracy across task sizes; (3) the task complexity hierarchy between SelectOne and SelectAll nearly disappears for all LLMs except Vicua, implying that it may be challenging for Vicuna to perform index selection tasks in general.
 Figure 7: Full average accuracy of 7 LLMs on the 4 classification-related tasks across the 6 classification benchmarks.
-Figure 9: Cumulative error density across positions for all benchmarks and LLMs across different task sizes.
-The task size “M/N” on the left side of the plots denotes the task size for the 3 text-pair benchmarks (i.e., MRPC, SNLI, and WiC) and for the 3 single-text benchmarks (i.e., SST-2, CoLA, AGNews), respectively.
-Figure 10: Full results for the two simplified index selection tasks versus the original tasks based on AGNews.
 Figure 8: Average accuracy of GPT-3.5 and Llama-3 70B on the other 6 reasoning benchmarks with multiple single-source problems.
 We leave out results where the number of parsable outputs is less than 50, e.g., GPT-3.5 on StrategyQA at task size > 2.
-Task SelectOne + CoT SelectAll + CoT BatchClf
-Model
-Vicuna 23.0 25.8 53.2 57.6 68.7
-Mistral 7B 38.3 47.5 56.8 60.5 71.2
-Mixtral 8x7B 47.4 38.7 65.1 58.4 73.4
-Llama-3 8B 39.0 41.8 62.4 59.0 73.3
-Llama-3 70B 59.4 67.1 72.9 79.2 79.4
-GPT-3.5 45.5 47.5 66.7 66.3 71.9
-GPT-4 66.3 71.8 78.8 81.8 81.9
-Overall 45.5 48.6 65.1 66.1 74.3
-Table 6: Aggregate average accuracy of SelectOne and SelectAll with and without 1-shot-CoT for each LLM. BatchClf performance is also provided for comparisons.
-B.6 CoT Helps the Index Selection Tasks Slightly but the Gap Remains
-We use 1-shot-CoT to prompt LLMs to do BatchClf first and then perform the two index selection tasks on the 6 benchmarks with a fixed task size 10.
-Table 6 shows that while overall LLMs benefit from CoT for both SelectOne and SelectAll, the benefits are generally larger for SelectOne than SelectAll (3.1% versus 1.0% improvement) and not consistent across models with Mixtral 8x7B performing even worse with CoT.
-The full results provided in Table D also show mixed effects across benchmarks.
-Interestingly, however, the task complexity hierarchy among BatchClf, SelectOne, and SelectAll remains in virtually all cases.
-We leave to future studies the investigation of how CoT may further help with the two selection tasks.
-Appendix C MPP: Reasoning
-This section provides additional details for Section 4. The prompt details for the experiments can be found in Appendix E.
+Appendix C Reasoning-Related Results
+This section provides additional details for Section 4.4. The prompt details for the experiments can be found in Appendix E.
 C.1 Construction of Mixed-Source Prompts
@@
 The numbers in the two parentheses are accuracies when treating “uncertain” answers as wrong answers, instead of discarding them.
-Appendix D Additional Results for Section 5
-This section describes the results of the three pretrained base LLMs and three FLAN-T5 models on Coin Flips at task sizes 1 and 2 from Section 5, shown in Table 7.
+Appendix D Further Analyses
+D.1 Positional Errors under BatchClf
+Fig 9 shows the full results regarding the positional errors 7 LLMs make across benchmarks and task sizes.
+We note that (1) distribution of the positional errors becomes more random (or even) as the task size increases for all LLMs; (2) in most cases, the positional errors distribute nearly randomly, showing no evidence of obvious positional biases, if any; (3) some LLMs may display more severe positional biases on some benchmarks with a certain task size, such as GPT-3.5 on SST-2 with task size 50, but overall this is rare.
+Figure 9: Cumulative error density across positions for all benchmarks and LLMs across different task sizes.
+The task size “M/N” on the left side of the plots denotes the task size for the 3 text-pair benchmarks (i.e., MRPC, SNLI, and WiC) and for the 3 single-text benchmarks (i.e., SST-2, CoLA, AGNews), respectively.
+D.2 Exploring Model-level Factors that may Enable MPP
+This section describes the results of the three pretrained base LLMs and three FLAN-T5 models on Coin Flips at task sizes 1 and 2 from Section 5.3, shown in Table 6.
 Similar to Table 3, Table 6 shows that instruction-tuned models perform much better on multi-problem prompts and that FLAN-T5 models can barely handle multi-problem prompts.
@@
 It is not clear whether "flip" means "flip" or "flip"…{The last sentence repeated for three more times.}…
-Benchmark SST-2 CoLA AGNews MRPC SNLI WiC
-Model Task
-Vicuna SelectOne 38.5 22.2 16.8 15.8 9.9 34.5
-+ CoT 52.6 53.9 12.1 12.1 11.5 12.5
-SelectAll 66.7 65.7 42.2 67.4 28.2 48.8
-+ CoT 81.7 69.2 40.9 63.9 40.9 48.9
-BatchClf 89.1 73.2 70.4 66.7 59.3 53.4
-[…]
-+ CoT 66.5 55.1 44.8 59.4 29.5 29.5
-SelectAll 68.3 63.6 48.2 58.4 54.7 47.9
-+ CoT 74.4 68.2 57.0 68.5 47.6 47.3
-BatchClf 79.6 76.8 78.1 74.0 69.1 49.6
-[…]
-+ CoT 41.5 67.3 12.9 48.4 25.1 37.2
-SelectAll 85.3 68.4 62.9 65.9 56.2 52.0
-+ CoT 49.0 67.4 57.8 66.5 58.9 51.1
-BatchClf 80.2 80.9 71.6 70.8 73.3 63.7
-[…]
-+ CoT 85.0 48.9 22.1 50.7 22.3 22.0
-SelectAll 85.0 62.9 59.6 64.8 50.1 52.2
-+ CoT 85.8 66.5 45.3 64.5 39.8 52.3
-BatchClf 92.0 76.8 79.5 62.8 69.2 59.5
-[…]
-+ CoT 94.7 79.7 75.2 67.9 69.4 15.8
-SelectAll 87.3 79.6 76.2 71.7 68.6 53.9
-+ CoT 96.1 83.7 85.5 73.9 77.3 58.4
-BatchClf 96.4 82.8 86.8 73.5 75.2 61.7
-[…]
-+ CoT 80.2 71.7 57.9 26.8 26.5 22.1
-SelectAll 83.1 79.7 68.3 61.5 55.8 51.6
-+ CoT 89.5 77.5 58.8 70.2 50.5 51.2
-BatchClf 95.0 81.8 86.4 52.7 57.4 58.0
-[…]
-+ CoT 84.5 81.9 67.8 60.1 77.0 59.6
-SelectAll 92.4 84.0 79.2 73.5 80.8 62.8
-+ CoT 91.3 83.7 85.4 76.2 83.3 71.0
-BatchClf 96.0 84.0 86.7 72.6 83.8 68.4
-Overall SelectOne 68.6 58.0 48.7 37.7 27.6 32.7
-+ CoT 72.1 65.5 41.8 46.5 37.3 28.4
-SelectAll 81.2 72.0 62.4 66.2 56.3 52.7
-+ CoT 81.1 73.7 61.5 69.1 56.9 54.3
-BatchClf 89.8 79.5 79.9 67.6 69.6 59.2
-Table 8: Full 1-shot-CoT results for all benchmarks and LLMS. The task size is fixed at 10.
 Appendix E Full Prompts
 E.1 Prompt Templates for SingleClf, BatchClf, SelectOne, and SelectAll
-Tables 13 to 14 show the complete prompt templates for the four task types (i.e., SingleClf, BatchClf, SelectOne, and SelectAll) tailored for SST-2, CoLA, AGNews, MPRC, SNLI, and WiC, respectively.
+Tables 11 to 12 show the complete prompt templates for the four task types (i.e., SingleClf, BatchClf, SelectOne, and SelectAll) tailored for SST-2, CoLA, AGNews, MPRC, SNLI, and WiC, respectively.
 While there are differences in the exact wording of a prompt template for each task type across the 6 classification benchmarks, each prompt template type shares a similar underlying structure and can be easily applied to other classification benchmarks.
@@
 This indicates the overall limited effects of rewording and restructuring prompts.
-E.2 Prompt Templates for the Two Index Selection Tasks
-Table 15 presents the two prompt templates used for the two simplified index selection tasks based on AGNews.
-E.3 One-Shot-CoT Prompt Templates for the Two Index Selection Tasks
-To construct a 1-shot-CoT exemplar, we first randomly sampled 10 examples from each benchmark that do not appear in our test examples.
-We then put these 10 examples into the prompt template for SelectOne or SelectAll, followed by a constructed answer.
-The answer first explicitly classifies the 10 examples into the corresponding class labels, and then perform index selection based on the generated class labels.
-Table 16 shows the complete answer templates for SelectOne and SelectAll for each classification benchmark.
-E.4 Prompt Template for Multi-problem Prompts for Reasoning Problems
-The prompt template for a multi-problem prompt made up of reasoning problems is straightforward, as described in Section 4.
+E.2 Prompt Template for Multi-problem Prompts for Reasoning Problems
+The prompt template for a multi-problem prompt made up of reasoning problems is straightforward, as described in Section 3.3.2.
 Below is a simple example prompt made up of 2 reasoning problems from CommonsenseQA.
@@
 Table 12: Prompt templates for WiC.
-Task Prompt template
-SelectOne Simplified Go over the \$num lines of text below and list the index numbers of the lines that contain the word ‘\$category’ according to the following instructions: If none of the texts contain the word ‘\$category,’ write ‘None.’ If all the texts contain the word ‘\$category,’ write ‘All.’ Otherwise, provide the index numbers of the texts that contain the word ‘\$category’ each on a separate line.
-Texts, one per line:
-\$texts
-‘None,’ ‘All,’ or the index numbers of the texts that contain the word ‘\$category,’ one per line:
-SelectAll Simplified In this task, each line of text contains one of four words ’Business,’ ’Sports,’ ’World,’ and ’Sci/Tech.’
-Go over the \$num lines of text below and list the index numbers of the lines that contain each word according to the following instructions: If none of the texts contain a particular word, write ‘None.’ If all the texts contain a particular word, write ‘All.’ Otherwise, provide the index numbers of the texts that contain each word.
-Output your responses in JSON format with the following keys: ‘business,’ ‘sports,’ ‘world,’ and ‘sci/tech.’ A formatted example output is provided below.
-{‘business’: [None/All or index numbers of texts containing ‘Business’], ‘sports’: [None/All or index numbers of texts containing ‘Sports’], ‘world’: [None/All or index numbers of texts containing ‘World’], ‘sci/tech’: [None/All or index numbers of texts containing ‘Sci/Tech’]}
-Texts, one per line:
-\$texts
-JSON output:
-Table 15: Prompt templates for simplified SelectOne and SelectAll based on AGNews.
-For SelectOne Simplified, we used an earlier prompt template that asks LLMs to produce outputs line by line, instead of a JSON object.
-As illustrated earlier, we find the effect of the two output formats for SelectOne to be minimal.
-Therefore, having a JSON output should not make a meaningful difference.
-Benchmark Task Answer template
-SST-2 SelectOne To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
-From here, we see that lines {indices} show positive sentiment. Therefore, the answer in JSON format is as follows: {final answer}
-SST-2 SelectAll To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
-From here, we see that lines {indices} show positive sentiment, and lines {indices} show negative sentiment. Therefore, the answer in JSON format is as follows: {final answer}
-CoLA SelectOne To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
-From here, we see that texts in lines {indices} are unacceptable. Therefore, the answer in JSON format is as follows: {final answer}
-CoLA SelectAll To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
-From here, we see that texts in lines {indices} are acceptable and texts in lines {indices} are unacceptable. Therefore, the answer in JSON format is as follows: {final answer}
-AGNews SelectOne To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
-From here, we see that texts in lines {indices} can be classified as ‘Sports.’. Therefore, the answer in JSON format is as follows: {final answer}
-AGNews SelectAll To solve this task, let’s first classify the 10 lines of text above, one per line: {labels}
-From here, we see that texts in lines {indices} can be classified as ‘Business,’ texts in lines {indices} can be classified as ‘Sports,’ texts in lines {indices} can be classified as ‘World,’ and texts in lines {indices} are ‘Sci/Tech.’ Therefore, the answer in JSON format is as follows: {final answer}
-MRPC SelectOne To solve this task, let’s first determine if text A is a paraphrase of text B for the 10 lines of text above, one per line: {labels}
-From there, we see that text pairs in lines {indices} are paraphrases. Therefore, the answer in JSON format is as follows: {final answer}
-MRPC SelectAll To solve this task, let’s first determine if text A is a paraphrase of text B for the 10 lines of text above, one per line: {labels}
-From there, we see that text pairs in lines {indices} are paraphrases and text pairs in lines {indices} are not. Therefore, the answer in JSON format is as follows: {final answer}
-SNLI SelectOne To solve this task, let’s first determine the inference relation between the premise and the hypothesis for the 10 lines of text above, one per line: {labels}
-From there, we see that text pairs in lines {indices} contain entailment inference relation. Therefore, the answer in JSON format is as follows: {final answer}
-SNLI SelectAll To solve this task, let’s first determine the inference relation between the premise and the hypothesis for the 10 lines of text above, one per line: {labels}
-From there, we see that text pairs in lines {indices} contain entailment inference relation, text pairs in lines {indices} contain contradiction inference relation, and text pairs in lines {indices} contain neutral inference relation.
-Therefore, the answer in JSON format is as follows: {final answer}
-WiC SelectOne To solve this task, let’s first determine if the target word is used with consistent meanings in the two subsequent contexts for the num lines of text above, one per line: {labels}
-From there, we see that text pairs in lines {indices} contain use the target words with inconsistent meanings. Therefore, the answer in JSON format is as follows: {final answer}
-WiC SelectAll To solve this task, let’s first determine if the target word is used with consistent meanings in the two subsequent contexts for the num lines of text above, one per line: {labels}
-From there, we see that text pairs in lines {indices} contain use the target words with inconsistent meanings and text pairs in lines {indices} do not.
-Therefore, the answer in JSON format is as follows: {final answer}
-Table 16: Answer templates for the two selection tasks for each classification benchmark.
-{indices}: a list of indices seperated by comma.
-{final answer}: answer in a JSON format specified by the SelectOne or SelectAll prompt.